### v1

# Insight Over Sight? Exploring the Vision-Knowledge Conflicts in Multimodal LLMs

Abstract
This paper explores the problem of commonsense-level vision-knowledge conflict in Multimodal Large Language Models (MLLMs), where visual information contradicts model’s internal commonsense knowledge (see Figure 1). To study this issue, we introduce an automated pipeline, augmented with human-in-the-loop quality control, to establish a benchmark aimed at simulating and assessing the conflicts in MLLMs. Utilizing this pipeline, we have crafted a diagnostic benchmark comprising 374 original images and 1,122 high-quality question-answer (QA) pairs. This benchmark covers two types of conflict targets and three question difficulty levels, providing a thorough assessment tool. Through this benchmark, we evaluate the conflict-resolution capabilities of nine representative MLLMs across various model families and find a noticeable over-reliance on textual queries. Drawing on these findings, we propose a novel prompting strategy, “Focus-on-Vision” (FoV), which markedly enhances MLLMs’ ability to favor visual data over conflicting textual knowledge. Our detailed analysis and the newly proposed strategy significantly advance the understanding and mitigating of vision-knowledge conflicts in MLLMs. The data and code are made publicly available. [^1] https://github.com/xyliu-cs/ConflictVIS
Figure 1: Illustration of the vision-knowledge conflict, where the visual input contradicts MLLM’s inherent knowledge. The MLLM’s response over-relies on its inherent commonsense knowledge.
1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2022; Touvron et al., 2023; OpenAI, 2023a; Meta AI, 2024) have reshaped the landscape of deep learning for their comprehensive capabilities in language understanding, reasoning, and generation (Wei et al., 2022a; b; Chen et al., 2023). This evolution has paved the way for the emergence of Multimodal Large Language Models (MLLMs) (Li et al., 2022a; 2023b; Lyu et al., 2023; Dai et al., 2023; Liu et al., 2024d; Zhu et al., 2023; Bai et al., 2023; Liu et al., 2024b; Wang et al., 2024; OpenAI, 2023b; Liu et al., 2024c; Tong et al., 2024; Gemini Team, 2024; OpenAI, 2024b), which integrate a vision model with the LLM to process and respond to visual information. MLLMs like GPT-4o (OpenAI, 2024b) and LLaVA-NeXT (Liu et al., 2024c) have exhibited remarkable proficiency across various vision-language tasks such as image captioning (Chen et al., 2015), visual question answering (VQA) (Antol et al., 2015), and visual reasoning (Johnson et al., 2017; Yue et al., 2024).
However, the persistence of knowledge conflicts in LLMs remains a significant challenge for MLLMs, often cited as a primary cause of hallucinations in these systems (Liu et al., 2023a; Guan et al., 2024). Introducing visual information into MLLMs generates a novel form of discrepancy termed “vision-knowledge conflict”, where the visual data contradicts the model’s pre-existing parametric knowledge. Prior research has mainly focused on assessing and comprehending these conflicts through the lens of factual knowledge by generating counterfactual images via image editing techniques (Guan et al., 2024). Nonetheless, the aspect of vision-knowledge conflict related to commonsense knowledge, [^2] To illustrate, a factual knowledge example is “The Eiffel Tower is 330 meters tall”, while a commonsense knowledge example is “Babies cry when they are hungry”. which proves more challenging for LLMs and MLLMs due to its implicit and nuanced nature (Li et al., 2022b), remains under-investigated.
To address this gap, we propose an automated pipeline with human-in-the-loop to develop a benchmark for simulating and analyzing commonsense-level vision-knowledge conflicts in MLLMs. Illustrated in Figure 2, our pipeline consists of four key modules: knowledge component extraction, counter-commonsense query construction, image generation, and question-and-answer (QA) generation. This framework streamlines the generation of counter-commonsense inputs from scratch and is modularly designed to facilitate the addition of further conflict categories and QA formats in the future. We demonstrate the application of our framework by developing the ConflictVis benchmark built on the Open Mind Common Sense dataset (Singh et al., 2002). The benchmark focuses on the critical aspects of commonsense knowledge in the visual scene: Subject, Action, and Place. It consists of 374 original images with 1,122 high-quality QA pairs spanning two conflict targets and three question difficulty levels, all verified by human annotators.
Using the crafted benchmark, we evaluate nine representative MLLMs from five model families, providing insights into model behaviors, the causes of conflicts, and effective methods designs to mitigate the negative effects of conflicts. Notably, when facing knowledge conflicts, MLLMs tend to over-rely on their parametric knowledge for simpler questions. For instance, the top commercial model, Claude-3.5-Sonnet, exhibits a memorization ratio (MR) on parametric knowledge of 43.6% on Yes-No questions, substantially higher than results on more complex subjective questions; the largest open-source model evaluated, LLaVA-NeXT-34B, shows a MR of 26.7% on Yes-No questions, three times more than that on subjective questions. Our detailed analysis of the failure cases reveals that MLLMs generally focus more on the textual query than on the visual context. Drawing on these observations, we revisit several existing training-free methods (Leng et al., 2024; Liu et al., 2024f; Wei et al., 2022b) to enhance the impact of visual context in answer generation. Interestingly, although Chain-of-Thought prompting (Wei et al., 2022b) improves model’s reasoning abilities, it directs MLLMs to over-utilize the parametric knowledge in the reasoning process, leading to inaccurate conclusions or refusals. From this understanding, we propose Focus-on-Vision (FoV) prompting to instruct MLLMs to prioritize visual information, which markedly improves the performance.
Our main contributions are summarized as follows:
  • We introduce an innovative framework to automatically construct counter-commonsense benchmarks from the ground up. This platform allows the flexible definition of conflict categories and QA formats, facilitating the large-scale creation of conflict samples with minimal human effort.
  • We present ConflictVis , a pioneering diagnostic benchmark specifically designed to evaluate vision-knowledge conflicts at the commonsense level in MLLMs. The benchmark is meticulously validated by human experts to guarantee the quality of data.
  • We perform extensive experiments on nine representative MLLMs, uncovering critical insights into model performance and the underlying mechanisms of the conflict. Leveraging these insights, we propose a novel Focus-on-Vision (FoV) prompting as an effective solution to significantly minimize vision-knowledge conflicts in MLLMs.
2 ConflictVis Benchmark
Figure 2: An automated pipeline with human quality control to construct images and question-answer pairs that conflict with commonsense knowledge.
This section outlines the pipeline and the data utilized to construct the ConflictVis benchmark.
2.1 Automated Pipeline
Commonsense knowledge refers to the information generally accepted by the majority of people about everyday life, encompassing practical principles on how the world functions (Singh et al., 2002). Inspired by the widely-used image captioning dataset (Chen et al., 2015), we explore a specific type of commonsense knowledge encapsulated as a triplet $<s,a,p>$, where $s$ is the Subject, $a$ is the Action, and $p$ is the Place of the action or subject. For example, the statement “a baby (Subject) drinking water (Action) on the bed (Place)” illustrates this format. In this structure, the Subject outlines the main object’s appearance, the Action describes the primary activity and its interactions with other relevant objects, and the Place highlights the background objects and setting. This three-part framework efficiently captures the vital details that are accurately reflected in an image.
We then describe the strategies for creating images and corresponding QA pairs that contradict the commonsense knowledge based on an input corpus. In general, we construct triplets containing low co-occurring concepts to serve as the counter-commonsense queries for image generation and fill pre-defined templates to generate the QA pairs. The pipeline contains the following four steps:
1. Extract Knowledge Components from Input Corpus.
The process commences with the extraction of knowledge components from the corpus, which contains human commonsense knowledge in natural language. Drawing inspiration from Li et al. (2021) on the creation of a compositional generalization test set, we identify frequently occurring components in the corpus and enumerate all possible combinations to generate knowledge triplets. To this end, we employ a transformer model pipeline to extensively annotate the syntactic labels, including dependency (DEP), Part-of-Speech (POS), and Named Entity (NE). Table 3 in Appendix lists the linguistic rules to identify each component:
  • Subject Phrases: Our framework examines noun chunks, selecting those where the head noun functions as the nominal subject, indicated by dependency labels “nsubj” or “nsubjpass”.
  • Action Phrases: The process starts by identifying the verb token (“VERB”) and then examines its dependents for direct objects marked as “dobj”. Upon finding such objects, our framework includes any determiners or compound modifiers present to form the complete action phrase. In addition, verbs are converted to the "VBG" form to align with the common practice in image captions.
  • Place Phrases: Distinguishing place phrases from other prepositional phrases relies on more than just syntactic labels. Accordingly, the framework employs semantic role labeling techniques by utilizing a BERT-based model (Gardner et al., 2017) to annotate each word’s semantic role. It then extracts the phrase with a location tag (“ARGM-LOC") that contains 3 to 4 words.
To develop a high-quality benchmark, our framework filters the top subject ($N_{S}$), action ($N_{A}$), and place ($N_{P}$) phrases, respectively, based on the occurrence. It also removes named entities from the subject and place phrases and merges similar phrases (e.g., “a doctor” and “the doctor”) by keeping the most frequently occurring variant.
2. Construct Counter-commonsense Query.
We aim to construct scenes with a single anomalous concept (i.e., target) that seldom co-occurs with the other concepts (i.e., the context pair). For instance, the Action component “fixing a computer” and the context pair (“the baby”, “on the bed”) results in the triplet “the baby on the bed fixing a computer”, which contradicts commonsense knowledge, as it is highly improbable for a baby to fix a computer in reality. Specifically,
  • The context pair should exhibit a high co-occurrence level, acting as a common background.
  • The target should display a low co-occurrence with the given context, representing the anomaly.
To find the high co-occurrence context pairs, our framework first groups context pairs by their target category, i.e., ContextAction consists of (Subject, Place) pairs and ContextPlace comprises (Subject, Action) pairs. These groups help to develop the focus of the questions in the next section. We omit ContextSubject to avoid the ambiguity and potential ethical issues for judging a subject’s identity.
We demonstrate how to construct queries with counter-commonsense Actions, and a similar approach can be applied to Places. For each subject, we pair it with all the Place components to establish context candidates and select the contexts based on the co-occurrence level. We use Normalized Pointwise Mutual Information (NPMI) to evaluate the co-occurrence of the context pairs ${\bf C}=(C_{X},C_{Y})$:
$\text{PMI}(C_{X};C_{Y})\equiv\log_{2}\frac{P(C_{X},C_{Y})}{P(C_{X})P(C_{Y})},\qquad\text{NPMI}(C_{X};C_{Y})\equiv\frac{\text{PMI}(C_{X};C_{Y})}{-\log_{2}P(C_{X},C_{Y})}.$ (1)
A higher NPMI score indicates a more frequent co-occurrence of the two context components. For each subject, we select Top-$K$ places with highest NPMI scores to form context pairs.
For each selected context pair, we construct counter-commonsense triplets by selecting a target Action component that is not compatible with the context pair. Concretely, we also use the NPMI metric (i.e. $\text{NPMI}(T;{\bf C})$) to evaluate the co-occurrence of the target $T$ and context pair ${\bf C}$. For each context pair, we select Top-$M$ actions with lowest NPMI scores to form counter-commonsense queries.
To better match real-world knowledge, we leverage an LLM trained on large-scale web data to calculate the probability $P(\cdot)$. For example, to calculate the joint probabilities for multiple co-occurring components,our framework forms a concatenated expression of the concept phrases (e.g., “a baby on the bed” or “a baby on the bed fixing a computer”) and inputs it into a pre-trained LLM to obtain the generative probability. Utilizing an LLM here offers an advantage over frequency counting for addressing the issue of zero frequencies for rare concept compositions, which often arises with limited dataset size. Since large language models are trained on large-scale data to learn the statistical tendencies of human language, their generative probability provides an approximation of the likelihood of certain phrase appearing in reality. For the probability of each individual component, we use its relative frequency within its category (Lin, 1999). Following generation, a manual review is conducted here to ensure the data quality (See Appendix B for further details).
3. Generate Image.
[…]
Figure 3: Illustration of generating questions across three difficulty levels, from easy to hard based on the conflict target (Action in this example.)
4. Generate Questions.
As illustrated in Figure 3, with the constructed counter-commonsense triplets, the pipeline generates questions across three difficulty levels, from easy to hard:
  • Yes-No Question: For counter-commonsense Actions and Places, our framework utilizes templates “Is/are [Subject Place] [Action]?” and “Is/are [Subject Action] [Place]?”, respectively. The correct response to this type of question is typically “Yes”.
  • Multiple-Choice Question: For these questions, the templates include “What is/are [Subject] doing [Place]?” and “Where is/are [Subject] [Action]?” for counter-commonsense Action and Place, respectively. For a predetermined option count $m$, our framework selects ($m-1$) seemingly relevant but incorrect targets from the candidates. To this end, the framework partitions candidates into $m-1$ bins $\{B_{1},...,B_{m-1}\}$ based on their NPMI scores relative to the context pair. We randomly sample one candidate from each bin, together with the counter-commonsense target, to formulate the question’s multiple-choice options.
  • Subjective Question: The question text in the multiple-choice question is reused, but no list of options is included. To guide the model toward generating a brief and specific response, we add the constraints like “Answer with a single phrase”, with slight adjustment according to the specific prompt for different MLLMs.
2.2 ConflictVis Benchmark Construction
Table 1: Statistics of the constructed benchmark.
Target #Triplets #Images #QAs
Action 188 171 513
Place 156 203 609
[…]
We introduce how to use the automated pipeline to construct the ConflictVis benchmark. We utilize top 100K sentences from the Open Mind Common Sense (OMCS) dataset (Singh et al., 2002) as our knowledge base. Developed by the MIT Media Lab, OMCS is a rich collection of human commonsense knowledge, consisting of assertive sentences contributed by thousands of volunteers. From this dataset, we extract the 100 most frequent Subject phrases ($N_{S}=100$), along with 150 most frequent phrases for Actions and Places ($N_{A}=N_{P}=150$). Using feedback from the LLM Vicuna-1.5-13b, [^3] https://huggingface.co/lmsys/vicuna-13b-v1.5 we select the top 3 phrases with the highest NPMI scores to create context pairs for each subject ($K=3$). In addition, we choose the top 3 targets with the lowest NPMI scores to assemble the candidate set of triplets ($M=3$). After manually filtering out unexpected combinations, the remaining triplets are used for image generation. The low-quality images are again manually removed. The specific guidelines and results in the quality control process are elaborated in the Appendix B. Following this two-stage generation and filtering process, the statistics of the resulting benchmark are detailed in Table 1. In total, ConflictVis consists of 374 original images with 1,122 QA pairs spanning two conflict targets and three difficulty levels, all verified by human experts. Figure 10 in Appendix shows the instances for the subject “a musician”.
3 Experiment
3.1 Setup
Models
To explore the behavior of MLLMs when encountering commonsense knowledge conflicts, we perform a comprehensive evaluation on 9 MLLMs with 7 representative open-source MLLMs ranging from 8B to 34B, and 2 state-of-the-art commercial MLLMs. This evaluation covers the following five model series: LLaVA (8B, 13B, 34B) (Liu et al., 2024b; c), BLIP-2 (12.1B, 13B) (Li et al., 2023b; Dai et al., 2023), Qwen-VL (9.6B) (Bai et al., 2023), GPT-4o (OpenAI, 2024b) and Claude-3.5-Sonnet (Anthropic, 2024).
Evaluation
We use Accuracy and Memorization Ratio (MR) (Longpre et al., 2021) as the main evaluation metrics in our experiment. Both metrics require classifying the model’s responses into different categories. For Yes-No and Multiple Choice questions, where there is a unique answer, we use exact matching for classification. For Subjective questions, due to the open-ended nature and the complexity of the task (i.e., evaluating both textual and visual relevance and correctness), we rely on human annotators to perform the classification (Guidelines and details are included in Appendix C).
3.2 Sanity Test
Figure 4: Responses to textual inputs.
Do the instances in ConflictVis really conflict with MLLMs’ commonsense knowledge?
Before evaluating MLLMs’ performance on ConflictVis , we first verify whether the counter-commonsense cases in ConflictVis genuinely create conflicts with the models’ commonsense knowledge. To this end, we use the textual inputs in ConflictVis and query the model with the following prompt format:
Based on common sense, is it possible for [context] [target]?
where the context and target are filled with the specific phrases from the input. For example, the query for the instance in Figure 3 is “Based on common sense, is it possible for the baby on the bed fixing a computer?” The model’s responses are categorized as “negate” (indicating a conflict), “accept” (indicating no conflict), and “other” (where the model does not give a direct answer). The results, shown in Figure 4, demonstrate that the vast majority of cases in ConflictVis indeed present valid knowledge conflicts that can be identified by MLLMs, with conflict rates ranging from 80.5% for GPT-4o to 99.2% for LLaVA-1.5-13B.
Figure 5: Model uncertainty on traditional VQA dataset and on our benchmark.
Is ConflictVis a more challenging benchmark for MLLMs?
Unlike the traditional VQA benchmarks, where visual information typically aligns with the model knowledge, our ConflictVis presents conflicts between these two sources of knowledge, making it a more challenging benchmark for MLLMs. To empirically validate our claim, we conduct a comparative analysis of model uncertainty between our ConflictVis and the conventional VQA benchmark (Antol et al., 2015). A higher degree of uncertainty in model predictions on a particular benchmark indicates that the benchmark poses a greater challenge for the model to navigate. Figure 5 shows the results, where higher entropy value denotes more uncertainty. It is evident that the model’s predictions are more uncertain on ConflictVis than the standard VQA benchmark, highlighting the increased challenge posed by our ConflictVis benchmark.
3.3 Benchmarking MLLMs
(a) Yes-No Questions
(b) Multi-Choice Questions
(c) Subjective Questions
Figure 6: MLLMs’ response distributions on different question types. (Responses in Vision category are considered correct.)
In this section, we assess how well the MLLMs handle conflicts between vision information and parametric knowledge using our ConflictVis benchmark. Following Longpre et al. (2021), we compare model predictions (i.e., responses) with and without the inclusion of the image and then quantify the instances where model’s predictions are influenced by the provided image:
  • Align with its predictions without the given image (Knowledge, $P_{K}$): This outcome suggests that the model generates its answer based on its pre-existing parametric knowledge rather than the visual information presented.
  • Yield a correct answer that aligns with the visual information (Vision, $P_{V}$): This indicates that the model is capable of adapting to new information, adjusting its output to match the visual input.
  • Result in a different answer altogether (Other): This demonstrates that the model can modify its output based on visual input, though the result does not perfectly align with the visual information.
Figure 6 presents the results. Ideally, the model should output the Vision answer, supported by visual information, rather than the Knowledge answer derived from textual training, or any Other answers. However, both open-source and commercial MLLMs often revert to producing the Knowledge answer, ignoring the visual information, to varying degrees. In general, commercial MLLMs perform better than the open-source counterparts, particularly in handling subjective questions. Among the open-source MLLMs, LLaVA-NeXT-34B, being the largest model, delivers the best overall performance, achieving an average prediction accuracy of 84.6%. The top commercial model, GPT-4o, reaches an average prediction accuracy of 89.9%.
Figure 7: Memorization ratio (MR) of MLLMs, where higher MR values indicate greater model overstability on parametric knowledge.
MLLMs are more likely to overly rely on parametric knowledge for simpler questions.
One interesting observation from Figure 6 reveals that MLLMs exhibit poorer performance on straightforward Yes-No questions. Notably, the open-source BLIP-2-12B achieved a low accuracy of 39.3%, while the commercial Claude-3.5-Sonnet managed only 56.4%.
We leverage Memorization Ratio (MR) as a metric to empirically estimate the model’s overstability — it’s brittleness to changing information (Longpre et al., 2021):
$MR=\frac{P_{K}}{P_{K}+P_{V}}.$ (2)
Figure 7 illustrates the MR across different question types. Notably, the MR values for Yes-No questions are significantly higher than for the other two question types across all MLLMs, indicating that MLLMs tend to overly rely on parametric knowledge when answering simpler questions.
(a) Counter-commonsense Action
(b) Counter-commonsense Place
Figure 8: MLLMs’ response distributions on two distinct categories of conflict targets.
Counter-commonsense actions are more challenging for MLLMs.
Our ConflictVis evaluates two distinct types of conflict targets: counter-commonsense actions and places. As shown in Figure 8, the results show that counter-commonsense actions pose more of a challenge than places do. Specifically, the average prediction accuracy for counter-commonsense actions stands at 73.9%, significantly lower than the 85.2% recorded for places. Similarly, the MR for actions is markedly higher at 23.9%, compared to 13.4% for places. This discrepancy could be attributed to the richer visual context available for identifying places (e.g., “in a studio”) as opposed to the sparse cues for actions (e.g., “paddling the boat”).
3.4 Steering MLLMs Towards Vision
Figure 9: Input-output relevance in a failure case where visual information is underutilized.
Underutilization of visual information.
Our primary experiment demonstrates that when MLLMs encounter visual inputs that contradict their pre-existing knowledge, around 18% of the responses generated are not aligned with the visual context but rather reflect the model’s prior knowledge. To further understand this discrepancy, we analyze the overall input-output relevancy inside MLLMs (Stan et al., 2024), seeking insights that could guide improvement strategies. Figure 9 illustrates a common case in which the model LLaVA-1.5-13B failed to accurately respond to a subjective question, highlighting the challenge. The model pays more attention to the textual query than to the visual context. Accordingly, the attended tokens {“musician”, “doing”, “studio”} guide the model to generate the answer “playing guitar” based on the parametric knowledge.
Table 2: MLLMs’ accuracy under different improvement methods.
Model Y-N MC Subj. Avg.
Qwen-VL-Chat-9.6B 69.8 80.5 89.3 79.9
[…]
     + FoV Prompting (this work) 82.4 83.2 87.4 84.3
LLaVA-1.5-13B 70.6 88.0 82.9 80.5
[…]
     + Pay Attention to Image (Liu et al., 2024f) 85.6 88.8 86.1 86.8
[…]
     + FoV Prompting (this work) 82.9 89.0 81.8 84.6
LLaVA-NeXT-34B 73.3 92.5 88.0 84.6
[…]
     + FoV Prompting (this work) 85.8 92.5 89.8 89.4
GPT-4o 74.9 97.1 97.9 89.9
[…]
     + FoV Prompting (this work) 75.9 96.5 98.9 90.5
Effectiveness of Existing Methods.
Based on the above observation, one straightforward improvement strategy is to enhance the impact of visual context when generating answers. We consider several training-free approaches along this direction and propose a novel prompting method:
  • Visual Contrastive Decoding (Leng et al., 2024) contrasts the output distributions derived from original and distorted visuals to ensure the generated content closely adheres to the visual inputs.
  • Pay Attention to Image (Liu et al., 2024f) adaptively adjusts and amplifies the attention weights assigned to image tokens, thereby giving greater prominence to visual elements.
  • Chain-of-Thought (CoT) prompting (Wei et al., 2022b) enables complex reasoning capabilities through intermediate reasoning steps.
While Wei et al. (2022b) demonstrates that CoT can enhance the commonsense reasoning capabilities of LLMs, its applicability to MLLMs in leveraging visual elements under conflict scenarios may not be directly transferable. In response to this problem, we propose a simple and effective prompting technique named Focus-on-Vision (FoV) prompting to explicitly instruct MLLMs to prioritize visual information, which can be seamlessly incorporated into current MLLMs:
[textual query] Please focus on the visual information.
Table 2 shows the results. We adhere to the methodologies outlined in the original papers by implementing Visual Contrastive Decoding (VCD) and Pay Attention to Image (PAI) techniques on the Qwen-VL-Chat-9.6B and LLaVA-1.5-13B models. Both approaches significantly improve the models’ performance, aligning with the outcomes reported in the literature for enhancing MLLMs trustworthiness. Interestingly, we notice that CoT prompting diminishes the performance of the models. Upon examining the failure cases, we find that though CoT envokes the commonsense reasoning abilities of MLLMs, these models can over-rely on their parametric knowledge to draw the conclusion, which results in an increased number of inaccurate response or refusals. We present the corresponding failure cases in Appendix E.2. In contrast, our Focus on Vision (FoV) prompting technique addresses this issue by guiding MLLMs to prioritize visual information, which boosts performances across models.
4 Related Work
Knowledge Conflicts.
Knowledge conflicts in LLMs can be divided into three categories (Xu et al., 2024b): within the retrieved context (Chen et al., 2022), within model’s parametric knowledge (Huang et al., 2023), and between the context and the model’s parametric knowledge (Xie et al., 2024; Wu et al., 2024a). These conflicts can lead to incorrect or inconsistent responses, undermining the model’s trustworthiness (Xu et al., 2024b; Xie et al., 2024). In MLLMs, these conflict types expand to multimodal inputs, where conflicts can occur between the input image and the text instruction (Liu et al., 2024e; Han et al., 2024), or when the image contains counterfactual information that contradicts the model’s parametric knowledge (Liu et al., 2024e; Guan et al., 2024).
To evaluate the conflicts between visual information and the parametric knowledge in MLLMs, HallusionBench (Guan et al., 2024) manually edits collected informational graphics to create normal-counterfactual image pairs, and verifies model consistency with simple Yes-No questions. AutoHallusion (Wu et al., 2024b) introduces an automated approach to generate counterfactual scenarios by altering correlated objects in images, utilizing Yes-No questions to probe object existence and spatial relationships. In the context of commonsense knowledge, PhD (Liu et al., 2024e) generates counter-commonsense images through manual collection and synthesis, employing Subjective questions for assessment. Our proposed ConflictVis distinguishes itself by:
  • Incorporating a wider range of question types and conflict targets, including Yes-No, Multiple-Choice, Subjective questions and counter-commonsense actions and places. This approach allows for a more thorough evaluation of MLLMs. Our findings indicate that MLLMs tend to disregard visual information more frequently when responding to simpler questions and conflicting action is more challenging for MLLMs to resolve.
  • Automatically generating counter-commonsense images based on the co-occurrence of knowledge components and crafting questions around these counter-commonsense aspects, ensuring better image-question alignment. With the aid of advanced text-to-image models, our framework can generate instances tailored to specific needs.
Hallucination in MLLMs. Hallucination in MLLMs refers to cases where the model generates descriptions that conflict with the given visual context. MLLMs’ hallucinations are typically categorized based on the type of incorrect information, such as non-existent objects, incorrect object attributes, and inaccurate object relations (Liu et al., 2024a). Relevant research has mainly focused on two key areas: developing benchmarks and metrics to detect hallucinations (Rohrbach et al., 2018; Li et al., 2023c; Liu et al., 2023a; Wang et al., 2023) and proposing strategies to mitigate them (Liu et al., 2023a; Leng et al., 2024; Huang et al., 2024; Liu et al., 2024f).
In the context of knowledge conflicts, hallucination occurs when the model gives precedence to its intrinsic knowledge rather than the visual context (Liu et al., 2023a; Guan et al., 2024; Liu et al., 2024e; a). Our study explores this issue by simulating knowledge conflicts that MLLMs may face during inference, and evaluates the effectiveness of different strategies to mitigate hallucinations (Leng et al., 2024; Liu et al., 2024f) by enhancing the model’s fidelity to visual inputs. Furthermore, we introduce a novel prompting technique that explicitly instructs MLLMs to prioritize visual information, thereby significantly and consistently reducing the occurrence of hallucinations.
Benchmarks for MLLMs. Traditional vision-language benchmarks are designed to assess independent skills, including image captioning (Chen et al., 2015), visual grounding (Rohrbach et al., 2016), and visual question answering (Antol et al., 2015; Hudson & Manning, 2019). However, with the emergence of MLLMs, there is a growing need for more comprehensive and tailored benchmarks. The strong zero-shot abilities and advanced language generation skills exhibited by MLLMs can render traditional benchmarks insufficient, as they may not account for the diversity of responses or the full range of MLLM capabilities. To address these limitations, researchers have developed more complex and comprehensive benchmarks to assess MLLMs across a wider range of capabilities (Yue et al., 2024; Fu et al., 2023; Li et al., 2023a; Liu et al., 2023c). Meanwhile, diagnostic benchmarks have also been proposed to evaluate specific challenges or traits in MLLMs, such as hallucination (Li et al., 2023c; Liu et al., 2023a), social bias (Howard et al., 2024), and model safety (Liu et al., 2023b). ConflictVis is a pioneering analytical benchmark to introduce conflicting visual contexts against the model’s commonsense knowledge, allowing us to investigate model performance in the face of vision-knowledge conflicts.
5 Conclusion
This paper introduces an innovative approach to understand and mitigate the significant challenge of vision-knowledge conflicts in MLLMs, with a particular focus on commonsense knowledge. By developing an automated pipeline and the ConflictVis benchmark, we offer a tool for systematically evaluating and understanding the nuances of these conflicts. Our extensive experiments across nine representative MLLMs reveal a tendency toward over-reliance on parametric knowledge, especially in simpler queries, which underscores the importance of the visual context in these models’ decision-making processes. Building on these insights, we evaluate three improvement methods and design a new Focus-on-Vision prompting to reinforce the significance of visual input, thereby reducing the impact of conflicts. Our work paves the way for future research aimed at improving the robustness and reliability of MLLMs. Our goal is to ensure these models can effectively integrate and interpret multimodal data, mitigating internal knowledge inconsistencies.
Appendix A Benchmark Details
A.1 Linguistic rules to extract knowledge components
Table 3: Linguistic rules to identify different components in the knowledge triplet.
Category Main Requirement Example
Subject p in noun_chunks && p.root.dep in ["nsubj", "nsubjpass"] a baby, a chef
Action p[0].pos == "VERB" && p[-1].dep == "dobj" fixing a computer
Place p.srl = "ARGM-LOC" && 3 <= p.length <= 4 at the bookstore
A.2 Examples in ConflictVis
(a) a musician paddling the boat in a studio
(b) a musician chopping some onions in a studio
(c) a musician playing the piano on a freeway
(d) a musician playing the piano in an ambulance
Figure 10: Some examples for the subject “a musician” in the ConflictVis Benchmark.
Appendix B Human Quality control for constructing ConflictVis
This section outlines the quality control procedures within our automated pipeline. As shown in the previous Figure 2, there are two required quality control points: (1) after generating knowledge triplets, and (2) after generating images. In practice, to reduce the workload for annotators at the first point, we subdivide it into three steps: (i) after extracting knowledge components, (ii) after generating context pairs, and (iii) after generating knowledge triplets. Below, we detail the guidelines applied at each stage.
B.1 Human Annotators
Our annotation team consists of four Ph.D. students majored in computer science. All annotators are compensated for their contributions.
B.2 Filtering Knowledge Triplets
Filtering knowledge components.
This step removes semantically abstract phrases in the Subject, Action, and Place categories, which are difficult to visualize. Annotators classify each phrase based on its concreteness:
  • Concrete (1). The phrase represents tangible objects or actions.
  • Abstract (0). The Phrase represents abstract ideas or concepts.
Table 4 provides some labeling examples at this step:
Table 4: Labeling examples for knowledge components.
Category Label Phrases
Subject 0 the statement, the story, the stock market, words
Subject 1 a patient, kids, politicians, whales, a sailor
Action 0 understand the event, lose weight, take care
Action 1 take a shower, use a computer, play chess
Place 0 in the event, in the newspaper, in your life
Place 1 on the ground, on the sidewalk, by the fire
In this step, the annotators perform labeling starting from the top 1000 most frequent phrases from each category in descending order, and keep the first 100 Subject phrases and the first 150 Action and Place phrases that meet the requirements.
Filtering context pairs.
The target of this step is to remove the unsatisfying phrases that remain in the candidate list after the automatic pre-filtering. To this end, our annotators conduct a binary classification on the context pairs based on their commonness in the real world. Specifically, the annotators are asked to examine the context tuples without the knowledge of their co-occurrence scores and label them as:
  • Common (1). The context tuple depicts a typical scene in the real world.
  • Uncommon (0). The context tuple depicts an unusual scene in the real world.
To better illustrate, we provide some labeling examples from our human annotators in Table 5 below:
Table 5: Labeling examples for context pairs.
Category Label Phrases
(Subject, Action) 0 (the cat, walking the dog), (a butcher, playing cards)
(Subject, Action) 1 (the cat, climbing a tree), (a butcher, slaughtering a pig)
(Subject, Place) 0 (cows, on the beach), (a gardener, in the desert)
(Subject, Place) 1 (cows, on a farm), (a gardener, in a greenhouse)
In this step, our annotators thoroughly label candidate context pairs within the Top K range provided by the pipeline. Context pairs labeled as 1 are retained for the next stage of generation.
Filtering knowledge triplets.
Following the preliminary filtering steps, the workload for this phase is significantly reduced. The goal here is to eliminate any unexpected triplets that are not captured in earlier filtering. Without access to the exact co-occurrence score between the context and target phrases, annotators perform a binary classification based on the commonness of the triplet’s expression. Specifically, the annotators are asked to label each triplet in the candidate list as:
  • Common (0). The context tuple depicts a typical scene in the real world.
  • Uncommon (1). The context tuple depicts an unusual scene in the real world.
Again, we provide some labeling examples from our human annotators in Table 6 for a better illustration:
Table 6: Labeling examples for knowledge triplets.
Category Label Phrases
((Subject, Place), Action) 0 ((the waitress, at the bar), using a vcr)
((Subject, Place), Action) 1 ((the waitress, at the bar), hitting a deer)
((Subject, Action), Place) 0 ((politicians, drinking alcohol), at the theater)
((Subject, Action), Place) 1 ((politicians, drinking alcohol), in the oven)
In this step, our annotators label all the candidate context-target triplets in the Top M range provided by the pipeline, and triplets with label 1 are used for subsequent image generation.
B.3 Filtering Images
The objective of this step is to select high-quality images that align closely with the text prompt. Each annotator is provided with an image and its corresponding text query, and evaluates the image based on two criteria:
  • Alignment with text prompt (0 or 1). The image should contain the key objects, actions, and background described in the text. The main focus should clearly represent the scene without ambiguity or misinterpretation.
  • Image quality (0 or 1). The image should be clear, free from significant distortions, artifacts, or unnatural effects like warping, blurring, or pixelation.
With these two guidelines, the annotators extensively label all the generated images, and the images with a total score of 2 are kept in the final dataset.
Below, we demonstrate some sample images that are below our quality standard:
(a) Cows in a field chopping carrots.
Labels: A(0), Q(1)
(b) The teacher in a classroom paddling the boat. Labels: A(1), Q(0)
(c) The farmer pouring milk in an ambulance.
Labels: A(0), Q(1)
(d) The pharmacist serving customers on a tree.
Labels: A(1), Q(0)
Figure 11: Examples of rejected images and their labels. A: Alignment with text prompt, Q: Quality of the image.
In this stage, we generate about 830 images in total and only keep 374 images, resulting in an average filter rate around 55%.
Appendix C Guidelines for Manual Labeling Open-end Responses
To label the correctness of MLLMs responses for the open-ended questions, we design our evaluation criteria following Xu et al. (2024a). Specifically, the annotators are asked to determine the correctness of MLLM’s responses based on two criteria:
  • Relevancy (0 or 1): The content of the response is reflected in the image and addresses the focus of the question.
  • Responsiveness (0 or 1): The response directly answers the question as instructed.
Responses with a total score of 2 are considered correct for accuracy calculation.
To classify responses as either knowledge-based or vision-based, annotators grade the semantic closeness between the current response and the reference responses on a scale from 0 to 2:
  • 0: The answers are entirely unrelated.
  • 1: The answers share similar concepts but are not identical or synonymous.
  • 2: The answers are identical or synonymous.
The candidate response is assigned to the category with a higher score (e.g., Vision or Knowledge). If both scores are 0, the response is categorized as “Other”.
Appendix D Additional experiment results
D.1 Summary of MLLMs performance on ConflictVis and general benchmarks
Table 7: Accuracy of MLLMs on general multimodal benchmarks and our ConflictVis . “Text”: number of text tokens, “Image”: number of image-text pairs. MMB denotes MMBench developed by Liu et al. (2023c), M3U denotes MMMU by Yue et al. (2024).
MLLMs Pretrain Data General Our Benchmark
Name Size Text Image MMB M3U Y-N MC Subj. Avg.
Open-Source MLLMs
Qwen-VL 9.6B 2.4T 3.6B 32.2 52.9 82.4 85.3 73.5
Qwen-VL-Chat 61.8 35.9 69.8 80.5 89.3 79.9
[…]
InstructBLIP 13B 35.7 68.2 88.8 81.8 79.6
[…]
LLaVA-NeXT 8B 15.0T 0.4B 71.7 41.7 70.3 86.6 85.0 80.7
34B 81.1 51.1 73.3 92.5 88.0 84.6
Close-Source MLLMs
Claude-3.5-Sonnet 79.7 68.3 56.4 90.6 96.8 81.3
[…]
Appendix E Failure Cases
E.1 ConflictVis failure cases
Figure 12: Failure case on yes-no question.
Figure 13: Failure case on multiple-choice question.
Figure 14: Failure case on subjective question.
E.2 CoT failure cases
Figure 15: CoT failure mode 1: The MLLM responses according to its commonsense knowledge.
Figure 16: CoT failure mode 2: The MLLM asserts that answering the question is impossible.

### v2

@@
-Insight Over Sight? Exploring the Vision-Knowledge Conflicts in Multimodal LLMs
+Insight Over Sight: Exploring the Vision-Knowledge Conflicts in Multimodal LLMs
 Abstract
-This paper explores the problem of commonsense-level vision-knowledge conflict in Multimodal Large Language Models (MLLMs), where visual information contradicts model’s internal commonsense knowledge (see Figure 1).
-To study this issue, we introduce an automated pipeline, augmented with human-in-the-loop quality control, to establish a benchmark aimed at simulating and assessing the conflicts in MLLMs.
-Utilizing this pipeline, we have crafted a diagnostic benchmark comprising 374 original images and 1,122 high-quality question-answer (QA) pairs.
-This benchmark covers two types of conflict targets and three question difficulty levels, providing a thorough assessment tool.
-Through this benchmark, we evaluate the conflict-resolution capabilities of nine representative MLLMs across various model families and find a noticeable over-reliance on textual queries.
-Drawing on these findings, we propose a novel prompting strategy, “Focus-on-Vision” (FoV), which markedly enhances MLLMs’ ability to favor visual data over conflicting textual knowledge.
-Our detailed analysis and the newly proposed strategy significantly advance the understanding and mitigating of vision-knowledge conflicts in MLLMs.
-The data and code are made publicly available.
-[^1] https://github.com/xyliu-cs/ConflictVIS
+This paper explores the problem of commonsense level vision-knowledge conflict in Multimodal Large Language Models (MLLMs), where visual information contradicts model’s internal commonsense knowledge.
+To study this issue, we introduce an automated framework, augmented with human-in-the-loop quality control, to generate inputs designed to simulate and evaluate these conflicts in MLLMs.
+Using this framework, we have crafted a diagnostic benchmark consisting of 374 original images and 1,122 high-quality question-answer (QA) pairs.
+The benchmark covers two aspects of conflict and three question types, providing a thorough assessment tool.
+We apply this benchmark to assess the conflict-resolution capabilities of nine representative MLLMs from various model families.
+Our results indicate an evident over-reliance on parametric knowledge for approximately 20% of all queries, especially among Yes-No and action-related problems.
+Based on these findings, we evaluate the effectiveness of existing approaches to mitigating the conflicts and compare them to our “Focus-on-Vision” prompting strategy.
+Despite some improvement, the vision-knowledge conflict remains unresolved and can be further scaled through our data construction framework.
+Our proposed framework, benchmark, and analysis contribute to the understanding and mitigation of vision-knowledge conflicts in MLLMs.
+Insight Over Sight: Exploring the Vision-Knowledge Conflicts
+in Multimodal LLMs
+Xiaoyuan Liu1,2thanks:   Work was done when Xiaoyuan, Wenxuan, Youliang, and Jen-tse were interning at Tencent.
+Wenxuan Wang3    Youliang Yuan1,2    Jen-tse Huang4    Qiuzhi Liu2    Pinjia He1thanks:   Pinjia He is the corresponding author.
+Zhaopeng Tu2 1School of Data Science, The Chinese University of Hong Kong, Shenzhen 2Tencent    3Renmin University of China    4Johns Hopkins University 1xiaoyuanliu@link.cuhk.edu.cn     1hepinjia@cuhk.edu.cn     2zptu@tencent.com
+1 Introduction
 Figure 1: Illustration of the vision-knowledge conflict, where the visual input contradicts MLLM’s inherent knowledge. The MLLM’s response over-relies on its inherent commonsense knowledge.
-1 Introduction
-Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2022; Touvron et al., 2023; OpenAI, 2023a; Meta AI, 2024) have reshaped the landscape of deep learning for their comprehensive capabilities in language understanding, reasoning, and generation (Wei et al., 2022a; b; Chen et al., 2023).
-This evolution has paved the way for the emergence of Multimodal Large Language Models (MLLMs) (Li et al., 2022a; 2023b; Lyu et al., 2023; Dai et al., 2023; Liu et al., 2024d; Zhu et al., 2023; Bai et al., 2023; Liu et al., 2024b; Wang et al., 2024; OpenAI, 2023b; Liu et al., 2024c; Tong et al., 2024; Gemini Team, 2024; OpenAI, 2024b), which integrate a vision model with the LLM to process and respond to visual information.
-MLLMs like GPT-4o (OpenAI, 2024b) and LLaVA-NeXT (Liu et al., 2024c) have exhibited remarkable proficiency across various vision-language tasks such as image captioning (Chen et al., 2015), visual question answering (VQA) (Antol et al., 2015), and visual reasoning (Johnson et al., 2017; Yue et al., 2024).
-However, the persistence of knowledge conflicts in LLMs remains a significant challenge for MLLMs, often cited as a primary cause of hallucinations in these systems (Liu et al., 2023a; Guan et al., 2024).
-Introducing visual information into MLLMs generates a novel form of discrepancy termed “vision-knowledge conflict”, where the visual data contradicts the model’s pre-existing parametric knowledge.
-Prior research has mainly focused on assessing and comprehending these conflicts through the lens of factual knowledge by generating counterfactual images via image editing techniques (Guan et al., 2024).
-Nonetheless, the aspect of vision-knowledge conflict related to commonsense knowledge, [^2] To illustrate, a factual knowledge example is “The Eiffel Tower is 330 meters tall”, while a commonsense knowledge example is “Babies cry when they are hungry”.
-which proves more challenging for LLMs and MLLMs due to its implicit and nuanced nature (Li et al., 2022b), remains under-investigated.
-To address this gap, we propose an automated pipeline with human-in-the-loop to develop a benchmark for simulating and analyzing commonsense-level vision-knowledge conflicts in MLLMs.
-Illustrated in Figure 2, our pipeline consists of four key modules: knowledge component extraction, counter-commonsense query construction, image generation, and question-and-answer (QA) generation.
-This framework streamlines the generation of counter-commonsense inputs from scratch and is modularly designed to facilitate the addition of further conflict categories and QA formats in the future.
-We demonstrate the application of our framework by developing the ConflictVis benchmark built on the Open Mind Common Sense dataset (Singh et al., 2002).
-The benchmark focuses on the critical aspects of commonsense knowledge in the visual scene: Subject, Action, and Place.
-It consists of 374 original images with 1,122 high-quality QA pairs spanning two conflict targets and three question difficulty levels, all verified by human annotators.
-Using the crafted benchmark, we evaluate nine representative MLLMs from five model families, providing insights into model behaviors, the causes of conflicts, and effective methods designs to mitigate the negative effects of conflicts.
-Notably, when facing knowledge conflicts, MLLMs tend to over-rely on their parametric knowledge for simpler questions.
-For instance, the top commercial model, Claude-3.5-Sonnet, exhibits a memorization ratio (MR) on parametric knowledge of 43.6% on Yes-No questions, substantially higher than results on more complex subjective questions; the largest open-source model evaluated, LLaVA-NeXT-34B, shows a MR of 26.7% on Yes-No questions, three times more than that on subjective questions.
-Our detailed analysis of the failure cases reveals that MLLMs generally focus more on the textual query than on the visual context.
-Drawing on these observations, we revisit several existing training-free methods (Leng et al., 2024; Liu et al., 2024f; Wei et al., 2022b) to enhance the impact of visual context in answer generation.
-Interestingly, although Chain-of-Thought prompting (Wei et al., 2022b) improves model’s reasoning abilities, it directs MLLMs to over-utilize the parametric knowledge in the reasoning process, leading to inaccurate conclusions or refusals.
-From this understanding, we propose Focus-on-Vision (FoV) prompting to instruct MLLMs to prioritize visual information, which markedly improves the performance.
-Our main contributions are summarized as follows:
+Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2022; Touvron et al., 2023; OpenAI, 2023a; Meta AI, 2024) have reshaped the landscape of deep learning for their comprehensive capabilities in language understanding, reasoning, and generation (Wei et al., 2022a, b; Chen et al., 2023).
+This evolution has paved the way for the emergence of Multimodal Large Language Models (MLLMs) (Li et al., 2022, 2023b; Lyu et al., 2023; Dai et al., 2023; Liu et al., 2024d; Zhu et al., 2023; Bai et al., 2023; Liu et al., 2024b; OpenAI, 2023b; Liu et al., 2024c; Tong et al., 2024; Gemini Team, 2024; OpenAI, 2024b; Meta AI, 2025), which integrate a vision model with the LLM to process visual information.
+Modern MLLMs like GPT-4o (OpenAI, 2024b) and LLaVA-NeXT (Liu et al., 2024c) have exhibited remarkable proficiency across various vision-language tasks such as image captioning (Chen et al., 2015), visual question answering (VQA) (Antol et al., 2015), and visual reasoning (Johnson et al., 2017; Yue et al., 2024).
+However, the persistence of knowledge conflicts in LLMs remains a significant challenge for MLLMs.
+Introducing visual information into MLLMs generates a novel form of discrepancy, a phenomenon we term “vision-knowledge conflict”, in which the visual data contradicts the model’s pre-existing parametric knowledge.
+Previous studies have demonstrated that LLMs can exhibit complex behavior when confronted with knowledge conflicts, oscillating between rigid adherence to their parametric knowledge and excessive sensitivity to contextual cues (Xie et al., 2024).
+This behavior brings substantial risks, especially when external information is critical for decision-making.
+Given the growing need for trustworthiness, real-time accuracy, and robustness in MLLM systems, it is imperative to further investigate and resolve these vision-knowledge conflicts, which are believed to be a primary cause of hallucinations (Liu et al., 2023a; Guan et al., 2024).
+While recent research has simulated these conflicts by manually crafting counterfactual images (Guan et al., 2024), there remains significant room for improvement in conflict categorization, question diversity, and image naturalness.
+Additionally, manually crafted benchmarks are limited in sample size and scalability, underscoring the need for a more automated pipeline to enable broader analysis.
+To this end, we propose an automated framework with human-in-the-loop to develop a benchmark for simulating and analyzing vision-knowledge conflicts in MLLMs at the commonsense level  [^1] For example, a commonsense knowledge statement is “babies cry when they are hungry”, as opposed to a factual knowledge statement like “the Eiffel Tower is 330 meters tall.”.
+Illustrated in Figure 2, our framework consists of 3 key modules: (1) knowledge component extraction, (2) counter-commonsense query construction, and (3) image and question-answer (QA) pair synthesis.
+This framework streamlines the generation of counter-commonsense inputs from scratch and is modularly designed to facilitate the addition of new conflict categories and QA formats in the future.
+We demonstrate the application of our framework by developing the ConflictVis benchmark, focusing on the aspects of Subject, Action, and Place.
+The benchmark consists of 374 original images with 1,122 high-quality QA pairs spanning two conflict targets and three question types, all manually verified.
+Using the crafted benchmark, we evaluate nine representative MLLMs from five model families, providing insights into model behaviors, the causes of conflicts, and effective approaches to mitigate the negative effects of conflicts.
+Notably, when facing knowledge conflicts, MLLMs tend to over-rely on their parametric knowledge for the answer, especially among Yes-No questions and action-related problems.
+For instance, the top commercial model, Claude-3.5-Sonnet, exhibits a memorization ratio on parametric knowledge of 43.6% on Yes-No questions, substantially higher than results on more complex Open-Ended questions.
+Regarding conflict type, the average memorization ratio for counter-commonsense action problems is 23.8%, 10.4 percentage points higher than that for the place problems.
+Our detailed analysis of the failure cases reveals that MLLMs generally underutilize visual information, relying on parametric knowledge to infer the answer based on textual clues.
+Drawing on these observations, we assess several existing improvement methods to enhance the impact of visual context in answer generation.
+Interestingly, although Chain-of-Thought prompting improves the reasoning abilities, it guides MLLMs to utilize parametric knowledge more during rationalization, often resulting in contradictory conclusions or refusals.
+In response, we propose “Focus-on-Vision” (FoV) prompting to directly instruct MLLMs to prioritize visual information, which markedly improves the model’s performance.
+Despite advancements by various mitigation approaches, the vision-knowledge conflict remains a persistent challenge.
+Our main contributions are summarized below:
   • We introduce an innovative framework to automatically construct counter-commonsense benchmarks from the ground up.
-This platform allows the flexible definition of conflict categories and QA formats, facilitating the large-scale creation of conflict samples with minimal human effort.
-  • We present ConflictVis , a pioneering diagnostic benchmark specifically designed to evaluate vision-knowledge conflicts at the commonsense level in MLLMs.
+This framework allows the flexible definition of conflict categories and QA formats, facilitating the large-scale creation of conflict samples with minimal human effort.
+  • We present ConflictVis, a pioneering diagnostic benchmark specifically designed to evaluate the commonsense level vision-knowledge conflicts in MLLMs.
 The benchmark is meticulously validated by human experts to guarantee the quality of data.
-  • We perform extensive experiments on nine representative MLLMs, uncovering critical insights into model performance and the underlying mechanisms of the conflict.
-Leveraging these insights, we propose a novel Focus-on-Vision (FoV) prompting as an effective solution to significantly minimize vision-knowledge conflicts in MLLMs.
+  • We benchmark nine representative MLLMs and evaluate the effectiveness of several improvement methods in addressing conflicts.
+Through this analysis, we demonstrate the significance of the vision-knowledge conflict.
+2 Related Work
+Knowledge Conflicts.
+Knowledge conflicts in LLMs can be divided into three categories (Xu et al., 2024c): within the retrieved context (Chen et al., 2022), within model’s parametric knowledge (Huang et al., 2023), and between the context and the model’s parametric knowledge (Xie et al., 2024; Wu et al., 2024a; Su et al., 2024).
+These conflicts can lead to incorrect or inconsistent responses, undermining the model’s trustworthiness (Xu et al., 2024c; Xie et al., 2024).
+In MLLMs, these conflict types expand to multimodal inputs, where conflicts can occur between the input image and the text instruction (Liu et al., 2024e; Han et al., 2024), or when the image contains counterfactual information that contradicts the model’s parametric knowledge (Liu et al., 2024e; Guan et al., 2024).
+To evaluate the conflicts between visual information and the parametric knowledge in MLLMs, HallusionBench (Guan et al., 2024) consists of manually edited informational graphics, forming (normal, counterfactual) image pairs used to assess model consistency through Yes-No questions.
+AutoHallusion (Wu et al., 2024b) introduces an automated approach to generate counterfactual scenarios by altering correlated objects in images, utilizing Yes-No questions to probe object existence and spatial relationships.
+In the context of commonsense knowledge, PhD (Liu et al., 2024e) generates counter-commonsense images through manual collection and synthesis, employing short-answer questions for assessment.
+In contrast, our proposed benchmark features an automated framework and a broad array of question types and conflict targets, enabling a more scalable and comprehensive evaluation of MLLMs.
+Hallucination in MLLMs.
+Hallucination in MLLMs refers to the situation where the model generates descriptions that conflict with the given visual context.
+MLLMs’ hallucinations are typically categorized based on the type of incorrect information, such as non-existent objects, incorrect object attributes, and inaccurate object relations (Liu et al., 2024a).
+Relevant research has mainly focused on two key areas: developing benchmarks and metrics to detect hallucinations (Rohrbach et al., 2018; Li et al., 2023c; Liu et al., 2023a; Wang et al., 2023) and proposing strategies to mitigate them (Liu et al., 2023a; Leng et al., 2024; Huang et al., 2024; Liu et al., 2024f).
+In the context of knowledge conflicts, hallucination occurs when the model gives precedence to its intrinsic knowledge rather than the visual context (Liu et al., 2023a; Guan et al., 2024; Liu et al., 2024e, a).
+Our study explores this issue by simulating knowledge conflicts and evaluating the effectiveness of various hallucination-mitigation approaches in resolving the conflicts.
+Benchmarks for MLLMs.
+Traditional vision-language benchmarks are designed to assess independent skills, including image captioning (Chen et al., 2015), visual grounding (Rohrbach et al., 2016), and visual question answering (Antol et al., 2015; Hudson and Manning, 2019).
+However, with the emergence of MLLMs, there is a growing need for more comprehensive and tailored benchmarks.
+The strong zero-shot abilities and advanced language generation skills exhibited by MLLMs make traditional benchmarks insufficient, as they may not account for the diversity of responses or the full range of MLLM capabilities.
+To address these limitations, researchers have developed more complex benchmarks to evaluate MLLMs across a wider range of tasks (Yue et al., 2024; Fu et al., 2023; Li et al., 2023a; Liu et al., 2023c).
+Meanwhile, diagnostic benchmarks have also been built to evaluate specific challenges or traits in MLLMs, such as hallucination (Li et al., 2023c; Liu et al., 2023a), social bias (Howard et al., 2024), and model safety (Liu et al., 2023b).
+ConflictVis is a pioneering analytical benchmark that presents conflicting visual contexts to challenge the model’s commonsense knowledge, enabling deeper investigations into model performance in the presence of vision-knowledge conflicts.
 3 ConflictVis Benchmark
-Figure 2: An automated pipeline with human quality control to construct images and question-answer pairs that conflict with commonsense knowledge.
-This section outlines the pipeline and the data utilized to construct the ConflictVis benchmark.
-2.1 Automated Pipeline
+Figure 2: An automated framework with human quality control to construct images and question-answer pairs that conflict with commonsense knowledge.
+This section outlines the framework and the data utilized to construct the ConflictVis benchmark.
+3.1 Automated Framework
 Commonsense knowledge refers to the information generally accepted by the majority of people about everyday life, encompassing practical principles on how the world functions (Singh et al., 2002).
-Inspired by the widely-used image captioning dataset (Chen et al., 2015), we explore a specific type of commonsense knowledge encapsulated as a triplet $<s,a,p>$, where $s$ is the Subject, $a$ is the Action, and $p$ is the Place of the action or subject.
-For example, the statement “a baby (Subject) drinking water (Action) on the bed (Place)” illustrates this format.
+Based on the widely-used image captioning dataset (Chen et al., 2015), we explore a specific type of commonsense knowledge encapsulated as a triplet $\langle\bm{s},\bm{a},\bm{p}\rangle$, where $s$ is the Subject, $a$ is the Action, and $p$ is the Place of the action or subject.
+For example, the statement “the waitress (Subject) washing dishes (Action) in the kitchen (Place)” illustrates this format.
 In this structure, the Subject outlines the main object’s appearance, the Action describes the primary activity and its interactions with other relevant objects, and the Place highlights the background objects and setting.
 This three-part framework efficiently captures the vital details that are accurately reflected in an image.
-We then describe the strategies for creating images and corresponding QA pairs that contradict the commonsense knowledge based on an input corpus.
-In general, we construct triplets containing low co-occurring concepts to serve as the counter-commonsense queries for image generation and fill pre-defined templates to generate the QA pairs.
-The pipeline contains the following four steps:
-1. Extract Knowledge Components from Input Corpus.
-The process commences with the extraction of knowledge components from the corpus, which contains human commonsense knowledge in natural language.
+We next describe our approach for generating images and corresponding QA pairs that challenge the commonsense knowledge in MLLMs.
+In general, we construct triplets of low co-occurring concepts to serve as counter-commonsense queries for the multimodal input generation.
+Our framework includes three stages, as depicted in Figure 2.
+We detail each stage below.
+Extract Knowledge Components
 Drawing inspiration from Li et al.
-(2021) on the creation of a compositional generalization test set, we identify frequently occurring components in the corpus and enumerate all possible combinations to generate knowledge triplets.
-To this end, we employ a transformer model pipeline to extensively annotate the syntactic labels, including dependency (DEP), Part-of-Speech (POS), and Named Entity (NE).
-Table 3 in Appendix lists the linguistic rules to identify each component:
-  • Subject Phrases: Our framework examines noun chunks, selecting those where the head noun functions as the nominal subject, indicated by dependency labels “nsubj” or “nsubjpass”.
-  • Action Phrases: The process starts by identifying the verb token (“VERB”) and then examines its dependents for direct objects marked as “dobj”.
-Upon finding such objects, our framework includes any determiners or compound modifiers present to form the complete action phrase.
-In addition, verbs are converted to the "VBG" form to align with the common practice in image captions.
-  • Place Phrases: Distinguishing place phrases from other prepositional phrases relies on more than just syntactic labels.
-Accordingly, the framework employs semantic role labeling techniques by utilizing a BERT-based model (Gardner et al., 2017) to annotate each word’s semantic role.
-It then extracts the phrase with a location tag (“ARGM-LOC") that contains 3 to 4 words.
-To develop a high-quality benchmark, our framework filters the top subject ($N_{S}$), action ($N_{A}$), and place ($N_{P}$) phrases, respectively, based on the occurrence.
-It also removes named entities from the subject and place phrases and merges similar phrases (e.g., “a doctor” and “the doctor”) by keeping the most frequently occurring variant.
-2. Construct Counter-commonsense Query.
-We aim to construct scenes with a single anomalous concept (i.e., target) that seldom co-occurs with the other concepts (i.e., the context pair).
-For instance, the Action component “fixing a computer” and the context pair (“the baby”, “on the bed”) results in the triplet “the baby on the bed fixing a computer”, which contradicts commonsense knowledge, as it is highly improbable for a baby to fix a computer in reality.
-Specifically,
-  • The context pair should exhibit a high co-occurrence level, acting as a common background.
-  • The target should display a low co-occurrence with the given context, representing the anomaly.
-To find the high co-occurrence context pairs, our framework first groups context pairs by their target category, i.e., ContextAction consists of (Subject, Place) pairs and ContextPlace comprises (Subject, Action) pairs.
-These groups help to develop the focus of the questions in the next section.
-We omit ContextSubject to avoid the ambiguity and potential ethical issues for judging a subject’s identity.
-We demonstrate how to construct queries with counter-commonsense Actions, and a similar approach can be applied to Places.
-For each subject, we pair it with all the Place components to establish context candidates and select the contexts based on the co-occurrence level.
-We use Normalized Pointwise Mutual Information (NPMI) to evaluate the co-occurrence of the context pairs ${\bf C}=(C_{X},C_{Y})$:
-$\text{PMI}(C_{X};C_{Y})\equiv\log_{2}\frac{P(C_{X},C_{Y})}{P(C_{X})P(C_{Y})},\qquad\text{NPMI}(C_{X};C_{Y})\equiv\frac{\text{PMI}(C_{X};C_{Y})}{-\log_{2}P(C_{X},C_{Y})}.$ (1)
-A higher NPMI score indicates a more frequent co-occurrence of the two context components. For each subject, we select Top-$K$ places with highest NPMI scores to form context pairs.
-For each selected context pair, we construct counter-commonsense triplets by selecting a target Action component that is not compatible with the context pair.
-Concretely, we also use the NPMI metric (i.e.
-$\text{NPMI}(T;{\bf C})$) to evaluate the co-occurrence of the target $T$ and context pair ${\bf C}$.
-For each context pair, we select Top-$M$ actions with lowest NPMI scores to form counter-commonsense queries.
-To better match real-world knowledge, we leverage an LLM trained on large-scale web data to calculate the probability $P(\cdot)$.
-For example, to calculate the joint probabilities for multiple co-occurring components,our framework forms a concatenated expression of the concept phrases (e.g., “a baby on the bed” or “a baby on the bed fixing a computer”) and inputs it into a pre-trained LLM to obtain the generative probability.
-Utilizing an LLM here offers an advantage over frequency counting for addressing the issue of zero frequencies for rare concept compositions, which often arises with limited dataset size.
-Since large language models are trained on large-scale data to learn the statistical tendencies of human language, their generative probability provides an approximation of the likelihood of certain phrase appearing in reality.
-For the probability of each individual component, we use its relative frequency within its category (Lin, 1999).
-Following generation, a manual review is conducted here to ensure the data quality (See Appendix B for further details).
-3. Generate Image.
+(2021) on the creation of a compositional generalization test set, we identify Subject, Action, and Place phrases as the most frequent elements in the corpus and build compounds based on them.
+To this end, we first employ a transformer model pipeline Honnibal et al.
+(2020) to extensively annotate the syntactic labels, including dependency (DEP), Part-of-Speech (POS), and Named Entity (NE).
+Then, we extract Subject, Action, and Place phrases from the corpus based on the predefined linguistic rules (See Appendix A.1).
+To ensure the data quality, our framework selects the top $N$ phrases from each category based on the frequency.
+It further refines the data by removing named entities from Subject and Place phrases and consolidating similar expressions (e.g., “a doctor” and “the doctor”) by retaining the most frequently occurring variant.
+This process results in three curated phrase lists for subsequent processing.
+Construct Counter-commonsense Query
+We aim to construct scenes with one single anomalous component (i.e., target) that seldom co-occurs with the others (i.e., the context).
+This objective can be formalized into two key requirements: (1) The context components should exhibit a high co-occurrence level, acting as a common background, and (2) The target component should display a low co-occurrence with the given context, representing the anomaly.
+To identify strongly co-occurring context pairs, our framework first groups context components by the target category, e.g., ContextAction consists of (Subject, Place) pairs.
+These groupings help to develop the focus of the questions in the next section.
+We omit ContextSubject to prevent ambiguity and mitigate potential ethical concerns related to subject identity.
+Next, our framework enumerate all the context combinations within the each group and computes their Normalized Pointwise Mutual Information (NPMI) scores.
+$\text{PMI}(C_{X};C_{Y})\equiv\log_{2}\frac{P(C_{X},C_{Y})}{P(C_{X})P(C_{Y})}$ (1)
+$\text{NPMI}(C_{X};C_{Y})\equiv\frac{\text{PMI}(C_{X};C_{Y})}{-\log_{2}P(C_{X},C_{Y})}$ (2)
+A higher NPMI score indicates a stronger co-occurrence between the two components.
+For each context group, the framework selects Top-$K$ context pairs with the highest NPMI scores to form a candidate pool.
+For each context pair in the pool, the framework generates counter-commonsense triplets by selecting a target component that is unusual given the context.
+Concretely, the framework evaluates the co-occurrence between each target $T$ and the context pair ${\bf C}$ using the NPMI score (i.e.
+$\text{NPMI}(T;{\bf C})$).
+For each context pair, the Top-$M$ targets with the lowest NPMI scores are retained to construct counter-commonsense queries.
+To better align the model knowledge, the framework queries an LLM trained on large-scale web data to estimate the probability $P(\cdot)$, and a manual review is conducted after the query generation to ensure quality (See Appendix A.2 for more details).
+Generate Multimodal Inputs
+Building on the constructed queries, our framework generates three types of questions, Yes-No, Multiple-Choice, and Open-Ended, along with their corresponding answers.
+To achieve this, the framework employs predefined question templates and fills in the relevant components accordingly.
+Example question-answer pairs are provided in Table 1.
+To generate corresponding images, our framework concatenates the triplet components into a caption-like expression and employs a prompt template to query a text-to-image model, as illustrated in Figure 2.
+Following image generation, human annotators perform quality control by filtering out low-quality images, such as those that appear distorted or misaligned with the input prompt.
+Detailed image filtering guidelines, along with illustrative examples, are provided in the Appendix B.3.
+Question Content Answer
+Yes-No Is the waitress in the kitchen signing a bill? Yes
+Multiple-Choice Question: What is the waitress doing in the kitchen?
+Options: (A) washing dishes. (B) riding a bike. (C) starting a fire. (D) signing a bill.
+D
+Open-Ended What is the waitress doing in the kitchen? signing a bill
+Table 1: Example question-answer pairs generated by our framework.
+3.2 ConflictVis Benchmark Construction
 […]
-Figure 3: Illustration of generating questions across three difficulty levels, from easy to hard based on the conflict target (Action in this example.)
-4. Generate Questions.
-As illustrated in Figure 3, with the constructed counter-commonsense triplets, the pipeline generates questions across three difficulty levels, from easy to hard:
-  • Yes-No Question: For counter-commonsense Actions and Places, our framework utilizes templates “Is/are [Subject Place] [Action]?” and “Is/are [Subject Action] [Place]?”, respectively.
-The correct response to this type of question is typically “Yes”.
-  • Multiple-Choice Question: For these questions, the templates include “What is/are [Subject] doing [Place]?” and “Where is/are [Subject] [Action]?” for counter-commonsense Action and Place, respectively.
-For a predetermined option count $m$, our framework selects ($m-1$) seemingly relevant but incorrect targets from the candidates.
-To this end, the framework partitions candidates into $m-1$ bins $\{B_{1},...,B_{m-1}\}$ based on their NPMI scores relative to the context pair.
-We randomly sample one candidate from each bin, together with the counter-commonsense target, to formulate the question’s multiple-choice options.
-  • Subjective Question: The question text in the multiple-choice question is reused, but no list of options is included.
-To guide the model toward generating a brief and specific response, we add the constraints like “Answer with a single phrase”, with slight adjustment according to the specific prompt for different MLLMs.
-2.2 ConflictVis Benchmark Construction
-Table 1: Statistics of the constructed benchmark.
 Target #Triplets #Images #QAs
@@
 […]
-We introduce how to use the automated pipeline to construct the ConflictVis benchmark.
-We utilize top 100K sentences from the Open Mind Common Sense (OMCS) dataset (Singh et al., 2002) as our knowledge base.
-Developed by the MIT Media Lab, OMCS is a rich collection of human commonsense knowledge, consisting of assertive sentences contributed by thousands of volunteers.
-From this dataset, we extract the 100 most frequent Subject phrases ($N_{S}=100$), along with 150 most frequent phrases for Actions and Places ($N_{A}=N_{P}=150$).
-Using feedback from the LLM Vicuna-1.5-13b, [^3] https://huggingface.co/lmsys/vicuna-13b-v1.5 we select the top 3 phrases with the highest NPMI scores to create context pairs for each subject ($K=3$).
-In addition, we choose the top 3 targets with the lowest NPMI scores to assemble the candidate set of triplets ($M=3$).
-After manually filtering out unexpected combinations, the remaining triplets are used for image generation.
-The low-quality images are again manually removed.
-The specific guidelines and results in the quality control process are elaborated in the Appendix B.
-Following this two-stage generation and filtering process, the statistics of the resulting benchmark are detailed in Table 1.
-In total, ConflictVis consists of 374 original images with 1,122 QA pairs spanning two conflict targets and three difficulty levels, all verified by human experts.
-Figure 10 in Appendix shows the instances for the subject “a musician”.
+Table 2: Statistics of the constructed benchmark.
 4 Experiment
@@
 Models
-To explore the behavior of MLLMs when encountering commonsense knowledge conflicts, we perform a comprehensive evaluation on 9 MLLMs with 7 representative open-source MLLMs ranging from 8B to 34B, and 2 state-of-the-art commercial MLLMs.
-This evaluation covers the following five model series: LLaVA (8B, 13B, 34B) (Liu et al., 2024b; c), BLIP-2 (12.1B, 13B) (Li et al., 2023b; Dai et al., 2023), Qwen-VL (9.6B) (Bai et al., 2023), GPT-4o (OpenAI, 2024b) and Claude-3.5-Sonnet (Anthropic, 2024).
+To explore the behavior of MLLMs when encountering vision-knowledge conflicts, we perform a comprehensive evaluation on 9 MLLMs including 7 representative open-source MLLMs ranging from 8B to 34B, and 2 state-of-the-art commercial MLLMs.
+This evaluation covers the following five model series: LLaVA (8B, 13B, 34B) (Liu et al., 2024b, c), BLIP-2 (12.1B, 13B) (Li et al., 2023b; Dai et al., 2023), Qwen-VL (9.6B) (Bai et al., 2023), GPT-4o (OpenAI, 2024b) and Claude-3.5-Sonnet (Anthropic, 2024).
+The diversity of model architectures and parameter sizes helps to enhance the generalizability of experiment results.
 Evaluation
@@
 Both metrics require classifying the model’s responses into different categories.
-For Yes-No and Multiple Choice questions, where there is a unique answer, we use exact matching for classification.
-For Subjective questions, due to the open-ended nature and the complexity of the task (i.e., evaluating both textual and visual relevance and correctness), we rely on human annotators to perform the classification (Guidelines and details are included in Appendix C).
+For Yes-No and Multiple-Choice questions, where there is a unique answer, we use exact matching for classification.
+For Open-Ended questions, due to the complexity of the task (i.e., evaluating both textual and visual relevance and correctness) and the instability of LLM evaluation (Stureborg et al., 2024), we rely on human annotators to perform the classification.
+Figure 3: MLLM response distributions on different question types.
 4.2 Sanity Test
-Figure 4: Responses to textual inputs.
 Do the instances in ConflictVis really conflict with MLLMs’ commonsense knowledge?
@@
 where the context and target are filled with the specific phrases from the input.
-For example, the query for the instance in Figure 3 is “Based on common sense, is it possible for the baby on the bed fixing a computer?” The model’s responses are categorized as “negate” (indicating a conflict), “accept” (indicating no conflict), and “other” (where the model does not give a direct answer).
+For example, the query for the instance in Figure 2 is “Based on common sense, is it possible for the waitress in the kitchen signing a bill?” The model’s responses (“Yes”, “No”, and others) are categorized as “accept” (indicating no conflict), “negate” (indicating a conflict), and “other” (where the model does not give a direct answer).
 The results, shown in Figure 4, demonstrate that the vast majority of cases in ConflictVis indeed present valid knowledge conflicts that can be identified by MLLMs, with conflict rates ranging from 80.5% for GPT-4o to 99.2% for LLaVA-1.5-13B.
-Figure 5: Model uncertainty on traditional VQA dataset and on our benchmark.
-Is ConflictVis a more challenging benchmark for MLLMs?
-Unlike the traditional VQA benchmarks, where visual information typically aligns with the model knowledge, our ConflictVis presents conflicts between these two sources of knowledge, making it a more challenging benchmark for MLLMs.
-To empirically validate our claim, we conduct a comparative analysis of model uncertainty between our ConflictVis and the conventional VQA benchmark (Antol et al., 2015).
-A higher degree of uncertainty in model predictions on a particular benchmark indicates that the benchmark poses a greater challenge for the model to navigate.
-Figure 5 shows the results, where higher entropy value denotes more uncertainty.
-It is evident that the model’s predictions are more uncertain on ConflictVis than the standard VQA benchmark, highlighting the increased challenge posed by our ConflictVis benchmark.
+Figure 4: MLLM responses to sanity test inputs.
 4.3 Benchmarking MLLMs
-(a) Yes-No Questions
-(b) Multi-Choice Questions
-(c) Subjective Questions
-Figure 6: MLLMs’ response distributions on different question types. (Responses in Vision category are considered correct.)
-In this section, we assess how well the MLLMs handle conflicts between vision information and parametric knowledge using our ConflictVis benchmark.
+In this section, we assess how well the MLLMs handle conflicts between visual information and parametric knowledge using our ConflictVis benchmark.
 Following Longpre et al.
-(2021), we compare model predictions (i.e., responses) with and without the inclusion of the image and then quantify the instances where model’s predictions are influenced by the provided image:
-  • Align with its predictions without the given image (Knowledge, $P_{K}$): This outcome suggests that the model generates its answer based on its pre-existing parametric knowledge rather than the visual information presented.
-  • Yield a correct answer that aligns with the visual information (Vision, $P_{V}$): This indicates that the model is capable of adapting to new information, adjusting its output to match the visual input.
-  • Result in a different answer altogether (Other): This demonstrates that the model can modify its output based on visual input, though the result does not perfectly align with the visual information.
+(2021), we compare model predictions (i.e., answers) with and without the inclusion of the image input and quantify the extent to which the model’s predictions are influenced by the presence of the image.
+Specifically, we compare the predictions in the current experiment, where the image is provided, to those obtained without the image using the sanity test prompts, and classify the current predictions as:
+  • Aligned with the prediction without the given image (Knowledge, $P_{K}$): This outcome suggests that the model generates its answer mainly based on its parametric knowledge rather than the visual information presented.
+  • Aligned with the visual information (Vision, $P_{V}$): This indicates that the model is capable of adapting to new information, adjusting its output to match the visual input.
+Such outputs are categorized as the correct answer.
+  • A different answer altogether (Other): This demonstrates that the model can modify its output based on visual input, though the result does not perfectly align with the visual information.
 Figure 3 presents the results.
-Ideally, the model should output the Vision answer, supported by visual information, rather than the Knowledge answer derived from textual training, or any Other answers.
-However, both open-source and commercial MLLMs often revert to producing the Knowledge answer, ignoring the visual information, to varying degrees.
-In general, commercial MLLMs perform better than the open-source counterparts, particularly in handling subjective questions.
-Among the open-source MLLMs, LLaVA-NeXT-34B, being the largest model, delivers the best overall performance, achieving an average prediction accuracy of 84.6%.
-The top commercial model, GPT-4o, reaches an average prediction accuracy of 89.9%.
-Figure 7: Memorization ratio (MR) of MLLMs, where higher MR values indicate greater model overstability on parametric knowledge.
-MLLMs are more likely to overly rely on parametric knowledge for simpler questions.
+Ideally, the model should only output the Vision answer, supported by visual information, rather than the Knowledge answer derived from textual training, or any Other answers.
+However, both open-source and commercial MLLMs revert to producing the Knowledge answer, ignoring the visual information, to varying degrees.
+In general, commercial MLLMs perform better than the open-source counterparts, particularly in handling open-ended questions.
+Among the open-source models, LLaVA-NeXT-34B achieves the highest prediction accuracy at 84.6%, while the top commercial model, GPT-4o, reaches 89.9%.
+Moreover, our analysis reveals that MLLMs’ uncertainty in answer generation is significantly higher on our ConflictVis compared to the classic VQA benchmark, indicating a greater challenge to the model performance (See Appendix C.2).
+Figure 5: Memorization ratio of MLLMs on different question types.
+MLLMs are more likely to overly rely on parametric knowledge for Yes-No questions.
 One interesting observation from Figure 3 reveals that MLLMs exhibit poorer performance on straightforward Yes-No questions.
-Notably, the open-source BLIP-2-12B achieved a low accuracy of 39.3%, while the commercial Claude-3.5-Sonnet managed only 56.4%.
-We leverage Memorization Ratio (MR) as a metric to empirically estimate the model’s overstability — it’s brittleness to changing information (Longpre et al., 2021):
+Notably, the open-source BLIP-2-12B achieves a low accuracy of 39.3%, while the commercial Claude-3.5-Sonnet manages only 56.4%.
+We leverage Memorization Ratio (MR) (Longpre et al., 2021) as a metric to empirically estimate the the extent to which the model adheres to its parametric knowledge in the presence of conflicting information.
 $MR=\frac{P_{K}}{P_{K}+P_{V}}.$ (3)
 Figure 5 illustrates the MR across different question types.
-Notably, the MR values for Yes-No questions are significantly higher than for the other two question types across all MLLMs, indicating that MLLMs tend to overly rely on parametric knowledge when answering simpler questions.
-(a) Counter-commonsense Action
-(b) Counter-commonsense Place
-Figure 8: MLLMs’ response distributions on two distinct categories of conflict targets.
+Notably, the MR values for Yes-No questions are significantly higher than the other two question types across all MLLMs, indicating that MLLMs tend to overly rely on parametric knowledge when answering Yes-No questions.
+We hypothesize that this could be due to the format of Yes-No questions, which directly present counter-commonsense expressions (e.g., “Is the waitress in the kitchen signing a bill?”) to the MLLMs, while other question types have a less direct phrasing.
+This counter-commonsense query may discourage the model from thoroughly analyzing the visual input, causing it to immediately output a negation.
 Counter-commonsense actions are more challenging for MLLMs.
-Our ConflictVis evaluates two distinct types of conflict targets: counter-commonsense actions and places.
-As shown in Figure 8, the results show that counter-commonsense actions pose more of a challenge than places do.
-Specifically, the average prediction accuracy for counter-commonsense actions stands at 73.9%, significantly lower than the 85.2% recorded for places.
-Similarly, the MR for actions is markedly higher at 23.9%, compared to 13.4% for places.
-This discrepancy could be attributed to the richer visual context available for identifying places (e.g., “in a studio”) as opposed to the sparse cues for actions (e.g., “paddling the boat”).
-3.4 Steering MLLMs Towards Vision
-Figure 9: Input-output relevance in a failure case where visual information is underutilized.
-Underutilization of visual information.
+Our ConflictVis evaluates two distinct types of conflict targets: counter-commonsense actions (e.g., the waitress in the kitchen signing a bill) and places (e.g., a doctor conducting an experiment at the theater).
+As shown in Figure 6, counter-commonsense action problems pose more of a challenge than places do.
+Specifically, MLLMs consistently exhibit a lower accuracy on counter-commonsense action problems, with an average output accuracy of 73.9%, notably lower than the 85.2% recorded for place problems.
+Similarly, the average Memorization Ratio for action problems is markedly higher at 23.8%, compared to 13.4% for place problems.
+This discrepancy could be attributed to the richer visual context typically available for identifying places (e.g., numerous background elements suggesting the location “at the theater”), in contrast to the relatively sparse and fine-grained visual cues required to recognize actions (e.g., subtle hand posture for “signing a bill”).
+Detailed results for each model can be found in the Appendix C.4.
+Figure 6: MLLM accuracy on Counter-Commonsense (CCS) Action problems and CCS Place problems.
+4.4 Mitigating Vision-Knowledge Conflicts
+Figure 7: Failure case analysis of input-output relevancy. The visual information is underutilized.
+Underutilization of visual information
 Our primary experiment demonstrates that when MLLMs encounter visual inputs that contradict their pre-existing knowledge, around 18% of the responses generated are not aligned with the visual context but rather reflect the model’s prior knowledge.
-To further understand this discrepancy, we analyze the overall input-output relevancy inside MLLMs (Stan et al., 2024), seeking insights that could guide improvement strategies.
-Figure 9 illustrates a common case in which the model LLaVA-1.5-13B failed to accurately respond to a subjective question, highlighting the challenge.
-The model pays more attention to the textual query than to the visual context.
-Accordingly, the attended tokens {“musician”, “doing”, “studio”} guide the model to generate the answer “playing guitar” based on the parametric knowledge.
-Table 2: MLLMs’ accuracy under different improvement methods.
-Model Y-N MC Subj. Avg.
-Qwen-VL-Chat-9.6B 69.8 80.5 89.3 79.9
-[…]
-[…]
-     + FoV Prompting (this work) 82.4 83.2 87.4 84.3
+To further investigate this discrepancy, we leverage the tool proposed by Stan et al.
+(2024) to analyze input-output relevancy scores, which identify the most relevant parts of the input to the generated output based on model attention.
+Figure 7(a) illustrates a case in which the model LLaVA-1.5-13B fails to accurately answer an open-ended question, representing a common failure pattern.
+In general, the model assigns greater weight to the textual input than to the visual context, resulting in an inaccurate response.
+In this case, the most attended tokens, “musician”, “doing”, “street”}, guide the model to generate “playing guitar” based on its parametric knowledge, while the image tokens, denoted by the <image> symbol in the bar plot, play a less significant role in answer generation.
+The image-output relevancy maps in Figure 7(b) further indicate that the model may not effectively leverage visual information during answer generation.
+In such cases, the model is more susceptible to vision-knowledge conflicts, as the visual input is largely underutilized.
+(More cases can be found in Appendix D.1.)
+Effectiveness of Existing Methods
+Based on the above observations, a straightforward strategy for improvement is to enhance the influence of visual context during the answer generation process.
+We consider several improvement approaches:
+  • Visual Contrastive Decoding (VCD) (Leng et al., 2024) contrasts the output distributions derived from original and noisy visuals to ensure the generated content adheres to the visual inputs.
+  • Pay Attention to Image (PAI) (Liu et al., 2024f) adaptively adjusts and amplifies the attention weights assigned to image tokens, thereby giving greater prominence to visual elements.
+  • Vision-Centric Reasoning (VR) employs structured intermediate reasoning steps to thoroughly analyze visual inputs, achieved by Chain-of-Thought (CoT) prompting (Wei et al., 2022b) or supervised finetuning on the long CoT multimodal reasoning dataset (Xu et al., 2024b).
+Furthermore, we propose a direct prompting technique, termed Focus-on-Vision (FoV) prompting, which explicitly instructs MLLMs to prioritize visual information.
+This can be seamlessly integrated into current MLLMs using the following template: [textual query] Please focus on the visual information.
+Model YN MC OE Avg.
 LLaVA-1.5-13B 70.6 88.0 82.9 80.5
 […]
-     + Pay Attention to Image (Liu et al., 2024f) 85.6 88.8 86.1 86.8
+     + PAI 85.6 88.8 86.1 86.8
+     + VR (CoT) 38.0 89.8 76.7 68.2
+     + VR (SFT) 64.0 87.4 88.5 80.0
+     + FoV 82.9 89.0 81.8 84.6
+Qwen-VL-Chat 69.8 80.5 89.3 79.9
 […]
-     + FoV Prompting (this work) 82.9 89.0 81.8 84.6
+     + VR (CoT) 79.7 65.8 77.8 74.4
+     + VR (SFT) 69.3 87.4 88.0 81.6
+     + FoV 82.4 83.2 87.4 84.3
 LLaVA-NeXT-34B 73.3 92.5 88.0 84.6
 […]
-     + FoV Prompting (this work) 85.8 92.5 89.8 89.4
+     + FoV 85.8 92.5 89.8 89.4
 GPT-4o 74.9 97.1 97.9 89.9
 […]
-     + FoV Prompting (this work) 75.9 96.5 98.9 90.5
-Effectiveness of Existing Methods.
-Based on the above observation, one straightforward improvement strategy is to enhance the impact of visual context when generating answers.
-We consider several training-free approaches along this direction and propose a novel prompting method:
-  • Visual Contrastive Decoding (Leng et al., 2024) contrasts the output distributions derived from original and distorted visuals to ensure the generated content closely adheres to the visual inputs.
-  • Pay Attention to Image (Liu et al., 2024f) adaptively adjusts and amplifies the attention weights assigned to image tokens, thereby giving greater prominence to visual elements.
-  • Chain-of-Thought (CoT) prompting (Wei et al., 2022b) enables complex reasoning capabilities through intermediate reasoning steps.
-While Wei et al.
-(2022b) demonstrates that CoT can enhance the commonsense reasoning capabilities of LLMs, its applicability to MLLMs in leveraging visual elements under conflict scenarios may not be directly transferable.
-In response to this problem, we propose a simple and effective prompting technique named Focus-on-Vision (FoV) prompting to explicitly instruct MLLMs to prioritize visual information, which can be seamlessly incorporated into current MLLMs:
-[textual query] Please focus on the visual information.
-Table 2 shows the results.
-We adhere to the methodologies outlined in the original papers by implementing Visual Contrastive Decoding (VCD) and Pay Attention to Image (PAI) techniques on the Qwen-VL-Chat-9.6B and LLaVA-1.5-13B models.
-Both approaches significantly improve the models’ performance, aligning with the outcomes reported in the literature for enhancing MLLMs trustworthiness.
-Interestingly, we notice that CoT prompting diminishes the performance of the models.
-Upon examining the failure cases, we find that though CoT envokes the commonsense reasoning abilities of MLLMs, these models can over-rely on their parametric knowledge to draw the conclusion, which results in an increased number of inaccurate response or refusals.
-We present the corresponding failure cases in Appendix E.2.
-In contrast, our Focus on Vision (FoV) prompting technique addresses this issue by guiding MLLMs to prioritize visual information, which boosts performances across models.
-4 Related Work
-Knowledge Conflicts.
-Knowledge conflicts in LLMs can be divided into three categories (Xu et al., 2024b): within the retrieved context (Chen et al., 2022), within model’s parametric knowledge (Huang et al., 2023), and between the context and the model’s parametric knowledge (Xie et al., 2024; Wu et al., 2024a).
-These conflicts can lead to incorrect or inconsistent responses, undermining the model’s trustworthiness (Xu et al., 2024b; Xie et al., 2024).
-In MLLMs, these conflict types expand to multimodal inputs, where conflicts can occur between the input image and the text instruction (Liu et al., 2024e; Han et al., 2024), or when the image contains counterfactual information that contradicts the model’s parametric knowledge (Liu et al., 2024e; Guan et al., 2024).
-To evaluate the conflicts between visual information and the parametric knowledge in MLLMs, HallusionBench (Guan et al., 2024) manually edits collected informational graphics to create normal-counterfactual image pairs, and verifies model consistency with simple Yes-No questions.
-AutoHallusion (Wu et al., 2024b) introduces an automated approach to generate counterfactual scenarios by altering correlated objects in images, utilizing Yes-No questions to probe object existence and spatial relationships.
-In the context of commonsense knowledge, PhD (Liu et al., 2024e) generates counter-commonsense images through manual collection and synthesis, employing Subjective questions for assessment.
-Our proposed ConflictVis distinguishes itself by:
-  • Incorporating a wider range of question types and conflict targets, including Yes-No, Multiple-Choice, Subjective questions and counter-commonsense actions and places.
-This approach allows for a more thorough evaluation of MLLMs.
-Our findings indicate that MLLMs tend to disregard visual information more frequently when responding to simpler questions and conflicting action is more challenging for MLLMs to resolve.
-  • Automatically generating counter-commonsense images based on the co-occurrence of knowledge components and crafting questions around these counter-commonsense aspects, ensuring better image-question alignment.
-With the aid of advanced text-to-image models, our framework can generate instances tailored to specific needs.
-Hallucination in MLLMs.
-Hallucination in MLLMs refers to cases where the model generates descriptions that conflict with the given visual context.
-MLLMs’ hallucinations are typically categorized based on the type of incorrect information, such as non-existent objects, incorrect object attributes, and inaccurate object relations (Liu et al., 2024a).
-Relevant research has mainly focused on two key areas: developing benchmarks and metrics to detect hallucinations (Rohrbach et al., 2018; Li et al., 2023c; Liu et al., 2023a; Wang et al., 2023) and proposing strategies to mitigate them (Liu et al., 2023a; Leng et al., 2024; Huang et al., 2024; Liu et al., 2024f).
-In the context of knowledge conflicts, hallucination occurs when the model gives precedence to its intrinsic knowledge rather than the visual context (Liu et al., 2023a; Guan et al., 2024; Liu et al., 2024e; a).
-Our study explores this issue by simulating knowledge conflicts that MLLMs may face during inference, and evaluates the effectiveness of different strategies to mitigate hallucinations (Leng et al., 2024; Liu et al., 2024f) by enhancing the model’s fidelity to visual inputs.
-Furthermore, we introduce a novel prompting technique that explicitly instructs MLLMs to prioritize visual information, thereby significantly and consistently reducing the occurrence of hallucinations.
-Benchmarks for MLLMs.
-Traditional vision-language benchmarks are designed to assess independent skills, including image captioning (Chen et al., 2015), visual grounding (Rohrbach et al., 2016), and visual question answering (Antol et al., 2015; Hudson & Manning, 2019).
-However, with the emergence of MLLMs, there is a growing need for more comprehensive and tailored benchmarks.
-The strong zero-shot abilities and advanced language generation skills exhibited by MLLMs can render traditional benchmarks insufficient, as they may not account for the diversity of responses or the full range of MLLM capabilities.
-To address these limitations, researchers have developed more complex and comprehensive benchmarks to assess MLLMs across a wider range of capabilities (Yue et al., 2024; Fu et al., 2023; Li et al., 2023a; Liu et al., 2023c).
-Meanwhile, diagnostic benchmarks have also been proposed to evaluate specific challenges or traits in MLLMs, such as hallucination (Li et al., 2023c; Liu et al., 2023a), social bias (Howard et al., 2024), and model safety (Liu et al., 2023b).
-ConflictVis is a pioneering analytical benchmark to introduce conflicting visual contexts against the model’s commonsense knowledge, allowing us to investigate model performance in the face of vision-knowledge conflicts.
+     + FoV 75.9 96.5 98.9 90.5
+Table 3: MLLM accuracy under different improvement methods. YN: Yes-No, MC: Multiple-Choice, OE: Open-Ended.
+Table 3 presents the results of the evaluated improvement methods.
+For VCD and PAI, we follow the methodologies outlined in the original papers and apply each technique to the corresponding applicable MLLMs, and observe performance improvements for both methods.
+To implement Vision-Centric Reasoning (VR), we explore two approaches: (1) applying Chain-of-Thought (CoT) prompting, and (2) fine-tuning models on the multimodal reasoning dataset introduced by Xu et al.
+(2024b).
+However, the results show minimal performance gains from finetuning, and vanilla CoT prompting can even degrade model accuracy.
+Upon analyzing failure cases, we find that teaching MLLMs to reason in a straightforward, step-by-step manner can lead to increased reliance on textual knowledge at test time.
+For example, our case study (in Appendix D.2) suggests that during CoT reasoning, models tend to rationalize the inputs using their existing commonsense knowledge.
+However, counter-commonsense inputs are inherently difficult to rationalize or explain.
+As the output text accumulates, it may become increasingly difficult for the model to accurately interpret the visual input given its own generated explanations.
+This often results in self-contradictory answers or refusals to answer.
+In contrast, our Focus-on-Vision (FoV) prompting mitigates this issue by explicitly directing the model’s attention to the visual input and consistently boosts model performance.
+Nevertheless, none of the approaches can entirely resolve the vision-knowledge conflict, especially for open-source models.
+For instance, LLaVA-1.5-13B and Qwen-VL-Chat still exhibit notable error rates of 13.2% and 15.7%, respectively, even when equipped with the most effective improvement method.
+This highlights the profound challenge posed by vision-knowledge conflict.
 5 Conclusion
-This paper introduces an innovative approach to understand and mitigate the significant challenge of vision-knowledge conflicts in MLLMs, with a particular focus on commonsense knowledge.
-By developing an automated pipeline and the ConflictVis benchmark, we offer a tool for systematically evaluating and understanding the nuances of these conflicts.
-Our extensive experiments across nine representative MLLMs reveal a tendency toward over-reliance on parametric knowledge, especially in simpler queries, which underscores the importance of the visual context in these models’ decision-making processes.
-Building on these insights, we evaluate three improvement methods and design a new Focus-on-Vision prompting to reinforce the significance of visual input, thereby reducing the impact of conflicts.
-Our work paves the way for future research aimed at improving the robustness and reliability of MLLMs.
-Our goal is to ensure these models can effectively integrate and interpret multimodal data, mitigating internal knowledge inconsistencies.
-Appendix A Benchmark Details
+In this paper, we develop an automated framework and construct the ConflictVis benchmark to systematically evaluating and understanding the nuances of vision-knowledge conflicts in MLLMs at the commonsense level.
+Our experiments across nine representative MLLMs reveal a tendency toward over-reliance on parametric knowledge, especially among Yes-No and action-related questions.
+We further evaluate three improvement methods and design a new prompting technique to mitigate the conflicts, but the problem remains notable.
+Collectively, our insights lay the groundwork for future research to develop more reliable MLLMs.
+Acknowledgment
+This paper was supported by the Guangdong Basic and Applied Basic Research Foundation (No.
+2024A1515010145) and the Shenzhen Science and Technology Program (Shenzhen Key Laboratory Grant No.
+ZDSYS20230626091302006).
+Limitations
+In this study, we focused on evaluating the reliably of multimodal large language models (MLLMs) under vision-knowledge conflict.
+While our work provides insights into the model performance and effectiveness of various mitigation approaches, it’s important to acknowledge certain limitations: (1) Limited root cause analysis: while we conducted failure case analysis using relevancy maps, the fundamental causes of the unexpected behaviors exhibited by MLLMs under vision-knowledge conflicts remain underexplored.
+We hypothesize that these behaviors may stem from biases in the training data, where MLLMs are predominantly aligned with image-text pairs that conform to commonsense expectations encoded during text pretraining, with limited exposure to counter-commonsense scenarios.
+To further validate this hypothesis, future work could involve controlled experiments comparing models trained with and without curated datasets designed for counter-commonsense alignment, in order to investigate the resulting behavior differences.
+(2) Unique probability model: due to budget constraints, our framework relies on Vicuna-1.5-13B to estimate co-occurrence probabilities, which may introduce model-specific features into the benchmark construction.
+Although our quality control processes help mitigate this issue, future work can explore more balanced probability estimation approaches, such as averaging probabilities across multiple models.
+Ethics Statements
+While our benchmark is designed to be safe and useful for research purposes, the developed framework for generating counter-commonsense inputs warrants careful consideration regarding its potential applications.
+In certain contexts, the framework may produce images that deviate from conventional norms, and if misused, could result in cases that some users may perceive as harmful or offensive.
+To mitigate such risks, we recommend implementing precautionary measures, including the use of a carefully curated and pre-filtered input corpus, as well as incorporating a manual verification process.
+Additionally, we encourage the use of text-to-image generative models through their official platforms and in accordance with their terms of use, which can help effectively leverage built-in safety mechanisms and reduce the likelihood of generating inappropriate or harmful content.
+Appendix A Benchmark Construction Details
 A.1 Linguistic rules to extract knowledge components
-Table 3: Linguistic rules to identify different components in the knowledge triplet.
+To systematically extract key knowledge components from textual inputs, our framework categorizes phrases into three main types, Subject, Action, and Place, based on syntactic and semantic annotations.
+Specifically, the framework first utilizes the transformer pipeline en_core_web_trf from spaCy (Honnibal et al., 2020) to annotate the sentences, and then performs the extraction based on the following procedures:
+  • Subject Phrases: Our framework examines noun chunks, selecting those where the head noun functions as the nominal subject, indicated by dependency labels “nsubj” or “nsubjpass”.
+  • Action Phrases: Our framework identifies the verb token (“VERB”) and then examines its dependents for direct objects marked as “dobj”.
+Upon finding such objects, the framework includes any determiners or compound modifiers present to form the complete action phrase.
+Verbs are converted to the “VBG” form to align with the common practice in image captions.
+  • Place Phrases: Distinguishing place phrases from other prepositional phrases relies on more than just syntactic labels.
+Accordingly, the framework employs semantic role labeling techniques by using the model structured-prediction-srl from AllenNLP (Gardner et al., 2017) to annotate each word’s semantic role.
+The framework then extracts the phrases with a location tag (“ARGM-LOC") that contains 3 to 4 words.
+In Table 4, we summarize the main linguistic rules employed to identify specific components.
 Category Main Requirement Example
-Subject p in noun_chunks && p.root.dep in ["nsubj", "nsubjpass"] a baby, a chef
-Action p[0].pos == "VERB" && p[-1].dep == "dobj" fixing a computer
-Place p.srl = "ARGM-LOC" && 3 <= p.length <= 4 at the bookstore
-A.2 Examples in ConflictVis
-(a) a musician paddling the boat in a studio
-(b) a musician chopping some onions in a studio
-(c) a musician playing the piano on a freeway
-(d) a musician playing the piano in an ambulance
-Figure 10: Some examples for the subject “a musician” in the ConflictVis Benchmark.
-Appendix B Human Quality control for constructing ConflictVis
-This section outlines the quality control procedures within our automated pipeline.
-As shown in the previous Figure 2, there are two required quality control points: (1) after generating knowledge triplets, and (2) after generating images.
+Subject p in noun_chunks && p.root.dep in [‘‘nsubj’’, ‘‘nsubjpass’’] a baby, a chef
+Action p[0].pos == ‘‘VERB’’ && p[-1].dep == ‘‘dobj’’ fixing a computer
+Place p.srl = ‘‘ARGM-LOC’’ && 3 <= p.length <= 4 at the bookstore
+Table 4: Linguistic rules to extract different components in the triplet.
+A.2 Probability Calculation
+To calculate the joint probabilities for multiple co-occurring components (e.g.
+$P(C_{X},C_{Y})$ or $P(C_{X},C_{Y},T)$), our framework forms concatenated expressions of the concept phrases (e.g., “the waitress in the kitchen” or “the waitress in the kitchen signing a bill”) and inputs them into a pre-trained LLM to obtain the generative probability.
+Utilizing an LLM here offers an advantage over frequency counting for addressing the issue of zero frequencies for rare concept compositions, which often arises with limited dataset size.
+Since large language models are trained on large-scale data to learn the statistical tendencies of human language, their generative probability provides an approximation of the likelihood of certain phrase appearing in reality.
+For the probability of each individual component, we use its relative frequency within its category (Lin, 1999).
+A.3 Question Generation Details
+With the constructed counter-commonsense triplets, our framework generates three types of questions:
+  • Yes-No Question: For counter-commonsense Actions and Places, our framework utilizes templates “Is/are [Subject Place] [Action]?” and “Is/are [Subject Action] [Place]?”, respectively.
+The correct response to this type of question is typically “Yes”.
+  • Multiple-Choice Question: For these questions, the templates are “What is/are [Subject] doing [Place]?” and “Where is/are [Subject] [Action]?” for counter-commonsense Action and Place, respectively.
+For a predetermined option count $m$, our framework selects ($m-1$) seemingly relevant but incorrect targets from the candidates.
+To this end, the framework partitions candidates into $m-1$ bins $\{B_{1},...,B_{m-1}\}$ based on their NPMI scores to the context pair.
+We randomly sample one candidate from each bin, together with the counter-commonsense target, to formulate the question’s multiple-choice options.
+We randomly shuffle the option list to eliminate any positional bias.
+  • Open-Ended Question: The question text in the multiple-choice question is reused, but no list of options is included.
+To guide the model toward generating a brief and specific response, we add the constraints like “Answer with a single phrase”, with slight adjustment according to the specific prompt for different MLLMs.
+A.4 Image Generation Details
+[…]
+A.5 Examples in ConflictVis
+In Figure 8, we demonstrate some samples generated by our framework. The conflict target is highlighted in red.
+(a) the hunter in the forest taking a note
+(b) the bear in the woods mailing a letter
+(c) a chef making bread at the theater
+(d) the knight riding a horse in a locker room
+Figure 8: Example counter-commonsense queries and their corresponding images generated in ConflictVis benchmark.
+A.6 Licenses
+[…]
+Appendix B Human Quality Control
+This section outlines the quality control procedures within our automated framework.
+As shown in Figure 2, there are two required quality control points: (1) after generating knowledge triplets, and (2) after generating images.
 In practice, to reduce the workload for annotators at the first point, we subdivide it into three steps: (i) after extracting knowledge components, (ii) after generating context pairs, and (iii) after generating knowledge triplets.
-Below, we detail the guidelines applied at each stage.
+In this section, we first describe the quality control guidelines applied at each stage of the framework, followed by the principles used for annotating open-ended responses.
 B.1 Human Annotators
-Our annotation team consists of four Ph.D. students majored in computer science. All annotators are compensated for their contributions.
+Our annotation team comprises four Ph.D.
+students specializing in computer science, all of whom are proficient in English.
+All annotators were fairly compensated based on their total working hours.
+Prior to annotation, we conducted training sessions to inform them that their tasks might include offensive content or identifying information, and clarified that their contributions would be used solely for research purposes to study model behavior.
+The annotation results were based on agreement among annotators.
+To ensure the reliability and objectivity of the evaluation process, each annotator independently performed the annotation tasks.
+Final labels were determined through majority voting, with group discussions held to resolve any ties.
 B.2 Filtering Knowledge Triplets
-Filtering knowledge components.
+Filtering knowledge components
 This step removes semantically abstract phrases in the Subject, Action, and Place categories, which are difficult to visualize. Annotators classify each phrase based on its concreteness:
   • Concrete (1). The phrase represents tangible objects or actions.
-  • Abstract (0). The Phrase represents abstract ideas or concepts.
-Table 4 provides some labeling examples at this step:
-Table 4: Labeling examples for knowledge components.
+  • Abstract or Unsafe (0). The phrase represents abstract concepts or contains offensive words or identifying information.
 Category Label Phrases
@@
 Subject 1 a patient, kids, politicians, whales, a sailor
-Action 0 understand the event, lose weight, take care
+Action 0 understand the event, lose weight, have sex
 Action 1 take a shower, use a computer, play chess
@@
 Place 1 on the ground, on the sidewalk, by the fire
+Table 5: Labeling examples for knowledge components.
+Table 5 provides some labeling examples at this step.
 In this step, the annotators perform labeling starting from the top 1000 most frequent phrases from each category in descending order, and keep the first 100 Subject phrases and the first 150 Action and Place phrases that meet the requirements.
-Filtering context pairs.
+Filtering context pairs
 The target of this step is to remove the unsatisfying phrases that remain in the candidate list after the automatic pre-filtering.
@@
   • Uncommon (0). The context tuple depicts an unusual scene in the real world.
-To better illustrate, we provide some labeling examples from our human annotators in Table 5 below:
+To better illustrate, we provide some labeling examples from our human annotators in Table 6:
+Category Label Phrases
+(Subject, Action) 0 (the cat, walking the dog),
+(a butcher, playing cards)
+(Subject, Action) 1 (the cat, climbing a tree),
+(a butcher, slaughtering a pig)
+(Subject, Place) 0 (cows, on the beach),
+(a gardener, in the desert)
+(Subject, Place) 1 (cows, on a farm),
+(a gardener, in a greenhouse)
 Table 6: Labeling examples for context pairs.
-Category Label Phrases
-(Subject, Action) 0 (the cat, walking the dog), (a butcher, playing cards)
-(Subject, Action) 1 (the cat, climbing a tree), (a butcher, slaughtering a pig)
-(Subject, Place) 0 (cows, on the beach), (a gardener, in the desert)
-(Subject, Place) 1 (cows, on a farm), (a gardener, in a greenhouse)
-In this step, our annotators thoroughly label candidate context pairs within the Top K range provided by the pipeline. Context pairs labeled as 1 are retained for the next stage of generation.
-Filtering knowledge triplets.
-Following the preliminary filtering steps, the workload for this phase is significantly reduced.
+In this step, our annotators thoroughly label candidate context pairs within the Top K range provided by the framework. Context pairs labeled as 1 are retained for the next stage of generation.
+Filtering knowledge triplets
+After the previous filtering steps, the workload for this phase is significantly reduced.
 The goal here is to eliminate any unexpected triplets that are not captured in earlier filtering.
@@
   • Uncommon (1). The context tuple depicts an unusual scene in the real world.
-Again, we provide some labeling examples from our human annotators in Table 6 for a better illustration:
-Table 6: Labeling examples for knowledge triplets.
 Category Label Phrases
@@
 ((Subject, Action), Place) 1 ((politicians, drinking alcohol), in the oven)
-In this step, our annotators label all the candidate context-target triplets in the Top M range provided by the pipeline, and triplets with label 1 are used for subsequent image generation.
+Table 7: Labeling examples for knowledge triplets.
+Again, we provide some labeling examples from our human annotators in Table 7 for a better illustration.
+In this step, our annotators label all the candidate context-target triplets in the Top M range provided by the framework, and triplets with label 1 are used for subsequent image generation.
 B.3 Filtering Images
@@
   • Image quality (0 or 1). The image should be clear, free from significant distortions, artifacts, or unnatural effects like warping, blurring, or pixelation.
-With these two guidelines, the annotators extensively label all the generated images, and the images with a total score of 2 are kept in the final dataset.
-Below, we demonstrate some sample images that are below our quality standard:
-(a) Cows in a field chopping carrots.
-Labels: A(0), Q(1)
-(b) The teacher in a classroom paddling the boat. Labels: A(1), Q(0)
-(c) The farmer pouring milk in an ambulance.
-Labels: A(0), Q(1)
-(d) The pharmacist serving customers on a tree.
-Labels: A(1), Q(0)
+With these two guidelines, the annotators extensively label all the generated images, and the images with a total score of 2 are kept for the subsequent stages.
+In Figure 9, we demonstrate some sample images that are below our quality standard.
+In this stage, we generate about 830 images in total and keep the best 374 images for the benchmark construction.
 Figure 9: Examples of rejected images and their labels.A: Alignment with text prompt, Q: Quality of the image.
-In this stage, we generate about 830 images in total and only keep 374 images, resulting in an average filter rate around 55%.
-Appendix C Guidelines for Manual Labeling Open-end Responses
+B.4 Guidelines for Manual Labeling Open-Ended Responses
 To label the correctness of MLLMs responses for the open-ended questions, we design our evaluation criteria following Xu et al.
 (2024a).
-Specifically, the annotators are asked to determine the correctness of MLLM’s responses based on two criteria:
+Specifically, the annotators are asked to determine the correctness of MLLMs’ responses based on two criteria:
   • Relevancy (0 or 1): The content of the response is reflected in the image and addresses the focus of the question.
@@
   • 2: The answers are identical or synonymous.
-The candidate response is assigned to the category with a higher score (e.g., Vision or Knowledge). If both scores are 0, the response is categorized as “Other”.
-Appendix D Additional experiment results
-D.1 Summary of MLLMs performance on ConflictVis and general benchmarks
-Table 7: Accuracy of MLLMs on general multimodal benchmarks and our ConflictVis .
-“Text”: number of text tokens, “Image”: number of image-text pairs.
-MMB denotes MMBench developed by Liu et al.
-(2023c), M3U denotes MMMU by Yue et al.
-(2024).
-MLLMs Pretrain Data General Our Benchmark
-Name Size Text Image MMB M3U Y-N MC Subj. Avg.
-Open-Source MLLMs
-Qwen-VL 9.6B 2.4T 3.6B 32.2 52.9 82.4 85.3 73.5
-Qwen-VL-Chat 61.8 35.9 69.8 80.5 89.3 79.9
+The candidate response is assigned to the category with a higher score (e.g., Vision or Knowledge).
+For example, a response is classified under the Vision category if it receives a vision score of 2 and a knowledge score of 1.
+Responses with both scores equal to 0 are categorized as “Other”.
+B.5 Human Performance
+In Table 8, we present the results of human annotators’ performance on our ConflictVis benchmark.
+The reported results are averaged across three annotators.
+The results indicate that human can achieve a higher accuracy than MLLMs.
+Question Type Accuracy
+Yes-No (YN) 93.0%
+Multiple-Choice (MC) 99.7%
+Open-Ended (OE) 91.7%
+Table 8: Human Performance on ConflictVis .
+Appendix C Experiment Details
+C.1 MLLM Generation Configurations
+For both open-source and closed-source models, we use the default generation configurations (e.g., temperature) to align with the typical usage.
+Question prompts for different MLLMs are crafted based on the prompt templates used in each model’s original evaluation.
+Specifically, we use the synthesized question text from the framework as the base and append a corresponding postfix tailored to each question type and model family.
+In practice, for the LLaVA series, GPT-4o, and Claude-3.5-Sonnet, the prompts are as follows:
+Question Type Prompt Format
+Multiple-Choice {question text} Answer with the option’s letter from the given choices directly.
+Yes-No {question text} Answer Yes or No directly.
+Open-Ended {question text} Answer no more than 5 words.
+For the BLIP-2 model family, the prompt postfixes are “Answer:” for MC and YN questions and “Short Answer:” for OE questions.
+For Qwen-VL, the postfixes include “Choice Answer:” for MC questions, while both YN and OE questions use “Answer:”.
+By contrast, Qwen-VL-Chat uses “Answer with the option’s letter from the given choices directly.” for MC questions, “Answer Yes or No directly.” for YN questions, and “Answer with a single word or phrase.” for OE questions.
+C.2 Uncertainty Measurement
+Unlike the traditional VQA benchmarks, where visual information typically aligns with the model knowledge, ConflictVis presents conflicts between the two sources of knowledge, making it a more challenging benchmark for MLLMs.
+To empirically validate our claim, we conduct a comparative analysis of model uncertainty between our ConflictVis and the conventional VQA benchmark (Antol et al., 2015).
+A higher degree of uncertainty in model predictions on a particular benchmark indicates that the benchmark poses a greater challenge for the model to navigate.
+Figure 10 shows the results, where a higher entropy value denotes more uncertainty.
+It is evident that the model’s predictions are more uncertain on ConflictVis than the standard VQA benchmark, highlighting the increased challenge posed by ConflictVis.
+Figure 10: Model uncertainty on traditional VQA dataset and on ConflictVis benchmark.
+To measure the uncertainty in model’s responses, we rely on the entropy. The average entropy $H_{\text{avg}}$ for a language model’s response $y$ is defined as:
+$H_{\text{avg}}=-\frac{1}{N}\sum_{n=1}^{N}\sum_{v\in V}p_{n}(v)\log p_{n}(v)$
+where $N$ is the number of tokens in the response $y$, $V$ is the vocabulary and $p_{n}(v)$ is the softmax probability of token $v$ in the vocabulary at position $n$ in the response sequence.
+To ensure a fair comparison, we first download the original VQA v1 dataset and extract an equal number of instances (374) for each question type (e.g., Multiple-Choice, Yes-No, and Open-Ended).
+Using the crafted subsets, we conduct controlled experiments under identical generation configurations to evaluate MLLMs on VQA v1 and our benchmark.
+The final uncertainty for each model is computed as the average entropy across all instances.
+C.3 Extended Analysis on No-type Questions
+To mitigate the potential dataset bias favoring “Yes” answers, we extend the Yes-No questions in our benchmark by constructing additional “No”-type questions and conduct further experiments.
+Specifically, for each original Yes-No query, we create a negative counterpart by replacing the counter-commonsense target with the most frequently co-occurring component from the corresponding option list.
+For example, given an image of a waitress signing a bill in the kitchen and the original question “Is the waitress in the kitchen signing a bill?” (expected answer: “Yes”), we generate the counterpart question “Is the waitress in the kitchen washing dishes?” (expected answer: “No”).
+We then reevaluate the MLLMs on these negative queries, with results summarized in Table 9.
+To compute the Memorization Ratio (MR), we follow the procedures outlined in Section 4.2: we first obtain the model’s responses without the image input and identify those accepted by the model, then evaluate these cases using the No-type question inputs, and finally compute MR according to Eq.
+3.
+The results indicate that these No-type questions still trigger a substantial number of failures across models, consistent with our original findings.
+Model Acc. (↑) MR (↓)
+LLaVA-NeXT-8B 81.3 21.8
+Qwen-VL-9.6B 93.0 11.2
+Qwen-VL-Chat-9.6B 71.7 31.5
+BLIP-2-12B 74.9 18.8
+InstructBLIP-13B 85.0 19.0
+LLaVA-1.5-13B 77.5 26.5
+LLaVA-NeXT-34B 82.1 20.8
+Claude-3.5-Sonnet 98.9 1.3
+GPT-4o 95.7 4.6
+Table 9: MLLM accuracy (higher is better) and Memorization Ratio (MR) (lower is better) on No-type Yes-No questions.
+C.4 Detailed Results on Conflict Targets
+As outlined in Section 4.3, here we provide detailed experimental results on counter-commonsense actions and places, as presented in Figure 11.
+Figure 11: MLLM response distributions on two distinct categories of conflict targets.
+C.5 Improvement Methods Configurations
+We present the detailed configurations for the improvement methods in Table 10.
+Method Configuration Method Configuration
 […]
-InstructBLIP 13B 35.7 68.2 88.8 81.8 79.6
 […]
-LLaVA-NeXT 8B 15.0T 0.4B 71.7 41.7 70.3 86.6 85.0 80.7
-34B 81.1 51.1 73.3 92.5 88.0 84.6
-Close-Source MLLMs
-Claude-3.5-Sonnet 79.7 68.3 56.4 90.6 96.8 81.3
-[…]
-Appendix E Failure Cases
-E.1 ConflictVis failure cases
+Table 10: Configurations for the improvement methods.
+Appendix D Case Study
+D.1 Extended Error Analysis
+(a) Failure case: The priest fixing a computer on the battlefield.
+(b) Failure case: The cook serving food in a shower.
+(c) Correct case: An own eating an apple in an ambulance(van).
+(d) Correct case: A musician chopping(cutting) some onions in a studio.
+Figure 12: Case study of input-output relevancy. Image tokens are underutilized in failure cases (a)(b), but effectively used in correct cases (c)(d).
+In Section 4.4, we identified a common failure pattern in which image tokens are underutilized during answer generation.
+In Figure 12, we present four additional examples from LLaVA-1.5-13B, spanning both counter-commonsense action and place categories.
+The first two are failure cases, while the bottom two are correct cases.
+In the failure cases, image tokens are significantly underutilized: in the bar plot, the <image> token, representing the image token with the highest relevancy score, is often less related to the output tokens than other textual tokens from the question description.
+The relevancy heatmaps also display only a few weakly attended regions.
+In contrast, for the correct cases, the <image> token typically has the highest relevancy scores for visual-dependent output tokens, and the heatmaps exhibit more numerous and strongly attended areas.
+Together, these examples highlight that the underutilization of visual information is a key factor contributing to model failures under vision-knowledge conflict.
+To address this issue, a stronger alignment between the image and the generated output is needed, guiding the development of more targeted and vision-aware improvement methods, as explored in the subsequent section.
+D.2 Failure Cases
+In this section, we present representative failure cases from our ConflictVis benchmark, spanning various models and question types.
+Specifically, Figures 13, 14, and 15 showcase failure cases for Yes-No, Multiple-Choice, and Open-Ended questions, respectively, while Figures 16 and 17 illustrate exacerbated failure cases when employing Chain-of-Thought prompting.
 Figure 13: Failure case on yes-no question.
 Figure 14: Failure case on multiple-choice question.
-Figure 14: Failure case on subjective question.
-E.2 CoT failure cases
-Figure 15: CoT failure mode 1: The MLLM responses according to its commonsense knowledge.
-Figure 16: CoT failure mode 2: The MLLM asserts that answering the question is impossible.
+Figure 15: Failure case on open-ended question.
+Figure 16: CoT failure mode 1: The MLLM reasons based on commonsense knowledge to derive an incorrect answer.
+Figure 17: CoT failure mode 2: The MLLM incorporates commonsense knowledge and asserts that the answer is impossible.